### v3

# Flavor diagonal nucleon charges using clover fermions on MILC HISQ ensembles

Abstract
We present lattice results for the flavor diagonal charges of the proton from the analysis of eight ensembles generated using 2+1+1-flavors of highly improved staggered quarks (HISQ) by the MILC collaboration. The calculation includes all the needed connected and disconnected contributions to nucleon three-point function. For extracting matrix elements using fits to the spectral decomposition of these correlation functions, two strategies to remove excited state contributions are employed and compared. To renormalize these charges, the 2+1-flavor mixing matrix is calculated in the RI-sMOM intermediate scheme on the lattice. The final results are presented in the $\overline{\text{MS}}$} scheme at scale 2 GeV. The axial charges for the proton are $g_{A}^{u}=0.781(25)$, $g_{A}^{d}=-0.440(39)$, and $g_{A}^{s}=-0.055(9)$; the tensor charges are $g_{T}^{u}=0.782(28)$, $g_{T}^{d}=-0.195(16)$, and $g_{T}^{s}=-0.0016(12)$; and the scalar charges are $g_{S}^{u}=9.39(88)$, $g_{S}^{d}=8.84(93)$, and $g_{S}^{s}=0.37(14)$. Results for the neutron are given by the $u\leftrightarrow d$ interchange. Results for the sigma terms are $\sigma_{\pi N}|_{\rm standard}=42(6)~{\rm MeV}$ from a “standard” analysis and $\sigma_{\pi N}|_{N\pi}=61(6)~{\rm MeV}$ from a $N\pi$” analysis that includes the contributions of multihadron $N\pi$ excited states as motivated by chiral perturbation theory. Our preferred value $\sigma_{\pi N}|_{N\pi}$ is consistent with the phenomenological extraction from $\pi-N$ scattering data. The strangeness content of the proton, for which the “standard” analysis is appropriate, is $\sigma_{s}|_{\rm standard}=35(13)~{\rm MeV}$.
preprint: LLNL-JRNL-858665, LA-UR-22053
I Introduction
Results for the flavor diagonal nucleon charges, $g_{A,S,T}^{u,d,s}$, are extracted from the matrix elements of axial, scalar, and tensor quark bilinear operators, ${\overline{q}}\Gamma q$ with the Dirac matrix $\Gamma=\gamma_{\mu}\gamma_{5},I,\sigma_{\mu\nu}$, respectively, within ground state nucleons. The lattice QCD calculations were done using Wilson-clover fermions on eight ensembles generated by the MILC collaboration with 2+1+1-flavors of highly improved staggered quarks (HISQ) [1]. The lattice parameters used in this clover-on-HISQ calculation are given in Tables 1 and 2.
The motivation for these calculations and much of the methodology used has already been published for $g_{A}^{q}$ in Ref. [2], $g_{T}^{q}$ in [3] and for the pion-nucleon sigma term, $\sigma_{\pi N}=m_{u,d}\times g_{S}^{u+d}$, in [4]. The charges $g_{A}^{u,d,s}$ give the contributions of the intrinsic spin of the quarks to the nucleon spin; the $g_{T}^{u,d,s}$ give the contribution of the quark electric dipole moment (EDM) operator to the nucleon EDM; and all three, $g_{A,S,T}^{u,d,s}$, give the coupling of dark matter or Higgs-like interactions with nucleons in the respective Lorentz channels. In addition, with $g_{S}^{u,d,s}$ in hand, we calculate the pion-nucleon sigma term, $\sigma_{\pi N}=m_{u,d}g_{S}^{u+d}$ and the strangeness content of the nucleon, $\sigma_{s}=m_{s}g_{S}^{s}$. We compare our results with those given in the latest FLAG report 2024 [5], a community review of calculations done until May 2024 by various lattice collaborations, and new results since then in Section VII.
This work supersedes our earlier publications [2, 3, 4] and updates in conference proceedings [6, 7]. It includes a number of improvements:
  • The calculation has been extended to eight ensembles described in Tables 1 and 2. The point $a06m220$ is new.
  • The disconnected contributions on all ensembles except $a12m220$ have now been calculated with operator insertion at all intermediate points $t$ between the nucleon source and the sink separated by Euclidean time $\tau$.
  • The statistics on most of the ensembles have been increased.
  • Results for $g_{S}^{u,d,s}$ are new. In Ref. [4], the result for only the renormalization group invariant pion-nucleon sigma term, $\sigma_{\pi N}=m_{u,d}\times g_{S}^{u+d}$, was presented. Using the fact that $m\overline{\psi}\psi$ is the mass term in the action, implies $Z_{m}\times Z_{S}=1$.
  • Complete results for the renormalization of quark bilinears for the clover-on-HISQ formulation, i.e., the 2+1-flavor mixing matrix, $Z_{A,S,T}$, are obtained nonperturbatively on the lattice using the regularization independent symmetric momentum subtraction (RI-sMOM) scheme [8, 9]. These $Z$s are then matched to the $\overline{MS}$ scheme and run to $2$ GeV using results from perturbation theory. Results in the basis (ud,u+d,s)(u-d,u+d,s), most relevant to this work, are given in Appendix B. Our previous calculations used $Z_{\text{isoscalar}}\approx Z_{\text{isovector}}$, i.e., $Z_{\text{u+d}}\approx Z_{\text{u-d}}$ for the renormalization constants of the axial and tensor operators. Having completed the full calculation, we find removing this approximation has not changed the results significantly because the corrections are small as can be inferred from the numbers in Appendix B.
  • Resolving and removing the contributions of excited states to nucleon correlation function continues to be a leading systematic. These artifacts have to be removed to get ground state matrix elements. In our recent lattice QCD calculation of the nucleon axial vector form factor $G_{A}(Q^{2})$ [10, 11, 12] and of the pion-nucleon sigma term $\sigma_{\pi N}$ [4], we presented evidence of larger-than-expected excited-state contributions (ESC) from $N\pi$ and $N\pi\pi$ multihadron excited states in these nucleon 3-point correlation functions. Motivated by these works, we study the impact of including $N\pi/N\pi\pi$ states in the analysis of all the flavor diagonal nucleon matrix elements.
    For correlation functions with resolvable signal of ESC, fits to get ground state matrix elements (GSME) now include two excited states in the spectral decomposition, and in each case compare fits with and without including the $N\pi$ excited state. These fits have been made using the full covariance matrix. To further control ESC, we vary $t_{\rm skip}$, the number of points skipped adjacent to the source and the sink, and the range of $\tau$ values included in the fits. For final results we take the average of these various fits, weighted by the Akaika Information Criteria (AIC) score [13].
  • The chiral-continuum-finite-volume (CCFV) extrapolation to the physical point, defined by $M_{\pi}\to 135$ MeV, $a\to 0$, and $M_{\pi}L\to\infty$, is done keeping the leading correction in each of the three variables.
This paper is organized as follows. A brief introduction to the lattice methods is given in Section II, and the strategies (models) used to remove excited state contributions are described in Section III. The averaging of models to get results for the charges using the AIC score is discussed in Section IV. The framework for obtaining the renormalized charges in the $\overline{\text{MS}}$} scheme at 2 GeV is described in Appendix B. The lattice calculation of the renormalization factors for the 2+1-flavor theory is done using the RI-sMOM scheme. The extrapolation of the renormalized data to the physical point using the CCFV fits is discussed in Section VI. A comparison with previous results and our conclusions are given in Section VII.
Ensemble ID $a$ (fm) $M_{\pi}$ (MeV) $\beta$ $C_{\rm SW}$ $am_{ud}$ $am_{s}$ $L^{3}\times T$ $M_{\pi}L$
$a15m310$ 0.1510(20) 320.6(4.3) 5.8 1.05094 -0.0893 -0.0210 $16^{3}\times 48$ 3.93
$a12m310$ 0.1207(11) 310.2(2.8) 6.0 1.05094 -0.0695 -0.018718 $24^{3}\times 64$ 4.55
$a12m220$ 0.1184(10) 227.9(1.9) 6.0 1.05091 -0.075 -0.02118 $32^{3}\times 64$ 4.38
$a09m310$ 0.0888(8) 313.0(2.8) 6.3 1.04243 -0.05138 -0.016075 $32^{3}\times 96$ 4.51
$a09m220$ 0.0872(7) 225.9(1.8) 6.3 1.04239 -0.0554 -0.01761 $48^{3}\times 96$ 4.79
$a09m130$ 0.0871(6) 138.1(1.0) 6.3 1.04239 -0.058 -0.0174 $64^{3}\times 96$ 3.90
$a06m310$ 0.0582(4) 319.6(2.2) 6.72 1.03493 -0.0398 -0.01841 $48^{3}\times 144$ 4.52
$a06m220$ 0.0578(4) 235.2(1.7) 6.72 1.03493 -0.04222 -0.01801 $64^{3}\times 144$ 4.41
Table 1: The lattice spacing $a$, the valence pion mass $M_{\pi}$, gauge coupling $\beta$, the Sheikholeslami-Wohlert coefficient $C_{SW}$ defining the clover term in the Wilson action, the light quark mass $am_{ud}=1/2\kappa_{ud}-4$, the strange quark mass $am_{s}=1/2\kappa_{s}-4$, the lattice volume $L^{3}\times T$, and the lattice size in units of $M_{\pi}$ for the eight ensembles analyzed for flavor diagonal charges $g_{\Gamma}^{q}$.
Ensemble ID $\sigma$ $\tau/a$ $N_{\text{conf}}^{\rm 2pt}$ $N_{\text{conf}}^{\rm conn}$ $N_{\text{LP}}$ $N_{\text{HP}}$ $N_{\text{conf}}^{l}$ $N_{\text{src}}^{l}$ $\frac{N_{\rm LP}^{l}}{N_{\rm HP}^{l}}$ $N_{\text{conf}}^{s}$ $N_{\text{src}}^{s}$ $\frac{N_{\rm LP}^{s}}{N_{\rm HP}^{s}}$
$a15m310$ 4.2 {6,7,8,9} 1917 1917 64 8 1917 2000 50 1917 2000 50
$a12m310$ 5.5 {8,10,12,14} 1013 1013 64 8 1013 10000 50 1013 8000 50
$a12m220$ 5.5 {8,10,12,14} 959 744 64 4 958(#) 11000 30 870 5000 50
$a09m310$ 7.0 {10,12,14,16} 2263 2263 64 4 1017 10000 50 1024 6000 50
$a09m220$ 7.0 {10,12,14,16} 964 964 128 8 712 8000 30 847 10000 50
$a09m130$ 7.0 {10,12,14,16} 1274 1290 128 4 1270 10000 50 994 10000/4000()($^{\dagger}$) 50
$a06m310$ 12.0 {18,20,22,24} 977 500 128 4 808 12000 50 976 10000/4000()($^{\dagger}$) 50
$a06m220$ 11.0/9.0 (*) {18,20,22,24} $10$ 649 64 4 1001 10000 50 1002 10000 50
Table 2: Parameters used in the calculation of the charges $g_{\Gamma}^{q}$ on the eight ensembles. The second column gives the smearing parameter $\sigma$ used in the creation of the covariant Gaussian smeared source for quark propagators. For ensemble $a06m220$, marked with (*), the connected contribution was done with $\sigma=11$ while the disconnected was done with $\sigma=9$. The list of source-sink time separations $\tau$ analyzed for the connected 3-point functions is given in column three. The disconnected contributions were calculated for all $\tau$ and all intermediate $t$ except on $a12m220$ for light quarks ($N^{l}_{\rm conf}$ marked with (#)). On this ensemble, data were collected with fixed $t_{\rm skip}=3$ and the four source time slices were fixed at $t_{src}=2,18,34,50$ rather than being chosen randomly. $N_{\text{conf}}^{\rm 2pt}$ and $N_{\text{conf}}^{\rm conn}$ give the number of configurations analyzed for 2-point and connected 3-point correlation functions, respectively. In the truncated solver with bias correction method, we used $N_{\text{LP}}$ low and $N_{\text{HP}}$ high precision measurements per configuration for the connected contributions. Columns 8–10 (11–13), specify the number of configurations, $N_{\text{conf}}^{l}$ ($N_{\text{conf}}^{s}$), analyzed for the light (strange) quark disconnected contributions with $N_{\rm src}^{l}$ ($N_{\rm src}^{s}$) $Z_{4}$ random noise sources and $N_{\rm LP}/N_{\rm HP}$ low to high precision measurements. In cases marked with $(^{\dagger})$, the $N^{s}_{\text{conf}}$ configurations were split into two sets and analyzed with different number of sources $N_{\rm src}^{s}$, and the results averaged. Results for isovector charges on full set of thirteen configurations including these has been published in Ref. [14].
II Lattice Methodology
Figure 1: The top row shows the connected (left) and disconnected (right) diagrams that contribute to the 3-point functions from which the matrix element of flavor-diagonal operators are extracted. The black and gray blobs denote the nucleon source and sink separated by Euclidean time $\tau$. The operator, shown by the symbol $\otimes$, is inserted at all intermediate times $t$ between the nucleon source and sink points. The bottom diagram (redraw of the top right) illustrates why the disconnected contribution for the scalar operator with $u$ and $d$ flavors can have an enhanced contribution due to the $N\pi$-intermediate state.
The charges $g_{A,S,T}^{u,d,s}$ are extracted from the forward matrix elements of the flavor diagonal quark bilinear operator, $\mathcal{O}_{\Gamma}^{q}(x)=\bar{q}(x)\Gamma q(x)$ with $q\in\{u,d,s\}$, inserted with zero momentum at time $t$ as illustrated in Fig. 1 using the relation
$\langle n(s_{f},{\mathbf{p}}=0)|{\bar{q}}\Gamma q|n(s_{i},{\mathbf{p}}=0)\rangle=g_{\Gamma}^{q}{\bar{u}}(s_{f})\Gamma u(s_{i})\,$, (1)
with $u(s)$ the nucleon spinor. These ME are obtained from the 3-point functions
$C_{\Gamma}^{\text{3pt}}(t;\tau)=\operatorname{Tr}[\mathcal{P}\langle 0|\mathcal{N}(\tau,\mathbf{p}=0)\mathcal{O}_{\Gamma}^{q}(t,\mathbf{q}=0)\bar{\mathcal{N}}(0)|0\rangle]$ (2)
where $\mathcal{P}$ is the spin projection defined in Ref. [15]. The interpolating operator $\mathcal{N}$} used to create/annihilate the nucleon at the source/sink time slices 0 and $\tau$ is
$\mathcal{N}(x)=\epsilon^{abc}\left[u^{aT}(x)C\gamma_{5}\frac{1\pm\gamma_{4}}{2}d^{b}(x)\right]u^{c}(x)\,.$ (3)
Here, $\{a,b,c\}$ are color indices and $C=\gamma_{4}\gamma_{2}$ is the charge conjugation matrix. The $\mathcal{N}$} is projected to zero momentum at the sink to get the forward matrix elements. Henceforth, we will drop the momentum indices as $\mathbf{p}=\mathbf{q}=0$ for all the calculations presented here. The goal is to design better $\mathcal{N}$}, and increase the statistical precision so that there is good signal at large enough $t$ and $\tau-t$ such that the extracted matrix elements are within the nucleon ground state. Otherwise ESC have to be removed using fits to the spectral decomposition of the correlation functions as discussed in Sec. III.
The nucleon 2- and 3-point correlation functions are calculated using the Wilson-clover fermion action with the Sheikholeslami-Wohlert coefficient $c_{SW}$ fixed to its tree-level tadpole improved value, $c_{SW}=1/u_{0}^{3}$ where $u_{0}$ is the tadpole factor given by the fourth root of the plaquette expectation value calculated on HISQ lattices that have been hypercubic (HYP) smeared [16].
Our calculation assumes isospin symmetry, i.e., $m_{u}=m_{d}\equiv m_{l}$, and the valence light quark mass $m_{l}$ is tuned so that clover-on-HISQ pion reproduces the HISQ-on-HISQ Goldstone (sea) pion mass $M_{\pi}^{\text{sea}}$. The valence strange quark mass is tuned so that the clover-on-HISQ pseudoscalar meson mass reproduces $M_{s\bar{s}}=\sqrt{m_{s}^{\text{sea}}/m_{l}^{\text{sea}}}M_{\pi}^{\text{ sea}}$ [15] where $m_{s}^{\text{sea}}/m_{l}^{\text{ sea}}$ is the ratio of HISQ (sea) quark masses and $M_{\pi}^{\text{sea}}$ is fixed at the physical value 135 MeV [1]. While this clover-on-HISQ mixed action formulation is not unitary and could have an addition systematic uncertainty, no obvious effects have been observed in our calculations, presumably because taking the continuum limit matching the Goldstone pseudoscalar spectra of the clover and HISQ formulations is sufficient.
The quark propagators used to construct the quark line diagrams shown in Fig. 1 are the inverse of the clover Dirac operator with a smeared (Wuppertal gauge covariant Gaussian [17]) source. The smearing parameter $\sigma$ is given in Table 2 and the same smearing is applied at the source and sink points. Efficacy of this smeared source in suppressing ESC was explored in Refs. [18, 19].
As illustrated in Fig. 1, the correlation function, $C_{\Gamma}^{\text{3pt}}(t;\tau)$ defined in Eq. (2), after Wick-contractions, is the sum of two types of quark lines diagrams called “connected” and “disconnected”:
$C_{\Gamma}^{\text{3pt}}(t;\tau)=C_{\Gamma}^{\text{conn}}(t;\tau)+C_{\Gamma}^{\text{disc}}(t;\tau).$ (4)
To reduce the computational cost, the connected diagrams are constructed using a coherent sequential source [20, 18] and the truncated solver with bias correction [21, 22, 18] methods. The quark loop with zero-momentum operator insertion, i.e., $\mathbf{q=0}$, for the disconnected contribution, $C_{\Gamma}^{\text{disc}}$, is estimated stochastically using $Z_{4}$ random noise sources. This calculation was accelerated with a combination of the truncated solver with bias correction method [15], and the hopping parameter expansion used as a pre-conditioner [23, 24, 15]. For the scalar case, the disconnected contribution is calculated using the vacuum subtracted operator $\mathcal{O}_{S}^{q}-\langle\mathcal{O}_{S}^{q}\rangle$ [25]. Additional details are given in Ref. [15]. The statistics for the connected and disconnected calculations are given in Table 2.
The calculation of the quark propagators and their contractions to construct correlation functions was done using the Chroma software suite [26]. On GPU nodes, we use the QUDA library [27, 28, 29] with the multigrid invertor [30, 31, 32] built into it [33].
In our previous works [2, 3], the removal of excited-state contamination (ESC) was carried out by making separate fits to $C^{\text{conn}}_{\Gamma}(t;\tau)$ and $C^{\text{disc}}_{\Gamma}(t;\tau)$. Similarly, the chiral-continuum (CC) extrapolations of $g_{\Gamma}^{q,\text{disc}}$ and $g_{\Gamma}^{q,\text{conn}}$ to $a\to 0$ and $M_{\pi}\to 135$ MeV were performed independently. This procedure may have introduced an unquantified systematic bias as discussed in Ref. [2]. This systematic has been removed in this work by making ESC fits to the sum and the subsequent CCFV extrapolation of the renormalized charges $g_{\Gamma}^{q}$. To understand the size of possible systematics in our previous work, we obtained results (ESC fits and CCFV extrapolation) both ways, and find the differences to be smaller than the statistical errors.
III Strategies to Remove Excited-State Contributions
The GSME $\langle 0|O_{\Gamma}^{q}|0\rangle$, and from them $g_{\Gamma}^{q;\text{bare}}$ using Eq. (1), are extracted from fits to the spectral decomposition of the spin-projected $C^{\textrm{3pt}}_{\Gamma}(t;\tau)$:
$C^{\textrm{3pt}}_{\Gamma}(t;\tau)$ $=\sum_{i,j=0}\mathcal{A}_{i}^{\ast}\mathcal{A}_{j}\langle i|O_{\Gamma}^{q}|j\rangle e^{-M_{i}t-M_{j}(t-\tau)}$ (5)
The challenge lies in removing all ESC since, even in our highest statistics ensembles, the right hand side of Eq. (5) is truncated at 3-states. While the ESC decrease exponentially with the mass gap $\Delta M$ and source-sink separation $\tau$ as $e^{-\Delta M\tau}$, the signal-to-noise ratio for $C^{\textrm{3pt}}_{\Gamma}$ degrades even faster exponentially as $e^{-(M_{N}-3/2M_{\pi})\tau}$ [34, 35]. The $C^{\textrm{3pt}}_{\Gamma}$ are well-measured only up to source-sink separation $\tau\approx 1.4\mathop{\rm fm}\nolimits$ and for $\tau\lesssim 1.4\mathop{\rm fm}\nolimits$, ESC are found to be significant. Three- state fits to remove these using Eq. (5) with $\mathcal{A}_{0}$, $M_{i}$ and the $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ left as free parameters are not stable with current statistics. Additional information on these parameters is needed.
To make robust fits to Eq. (5), the key parameters are the ground state amplitude $\mathcal{A}_{0}$ and the finite volume spectrum $M_{i}$. If these are known a priori, then to extract $\langle 0|O_{\Gamma}^{q}|0\rangle$ only the combinations $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ involving the excited states remain as free parameters. This then becomes a tractable problem. Note, these $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ are not used in any subsequent analysis.
In principle, the $\mathcal{A}_{0}$ and $M_{i}$ can be obtained from fits to the spectral decomposition of the 2-point function:
$C^{\textrm{2pt}}(\tau)=\sum_{i=0}^{3}|\mathcal{A}_{i}|^{2}e^{-M_{i}\tau}\,.$ (6)
To make n-state fits to $C^{\text{3pt}}_{\Gamma}(t;\tau)$ with input from $C^{\textrm{2pt}}(\tau)$, we need $C^{\textrm{2pt}}(\tau)$ truncated to $n+1$ (or higher) number of states presuming that the effects of all the neglected states are subsumed in the parameters of the $n+1$ (and higher) state. Of these, $M_{1}$ is the most important for controlling ESC in fits to $C^{\text{3pt}}_{\Gamma}(t;\tau)$ with the contamination falling exponentially with the mass gap. Unfortunately, with current statistics, 4-state fits to Eq. (6) give large regions of parameter space with roughly the same $\chi^{2}/dof$. We, therefore, resort to using physics motivated priors to determine the $M_{i}$.
The second challenge is that fits to $C^{\textrm{2pt}}(\tau)$ do not expose multihadron (such as $N\pi$, $N\pi\pi,\ldots$) excited states. The reason is that the coupling of multihadron states with nucleon quantum numbers to $\cal{N}$ is suppressed by factors of $1/L^{3}$ for each extra state [36]. On the other hand, as discussed later, their contribution to the matrix elements of certain operators is enhanced and compensates for the volume suppression. Therefore, in cases with enhanced contributions, these states need to be included in fits to $C^{\text{3pt}}_{\Gamma}(t;\tau)$ because the mass of $N(1)\pi(-1)$ (or $N(0)\pi(0)\pi(0)$) state is about 1230 MeV, which is significantly smaller than the first radial excitation $N(1440)$. In short, our first goal is to determine all the excited states, i.e., their $M_{i}$, that make significant contributions.
Our overall approach consists of making simultaneous fits to $C^{\textrm{2pt}}(\tau)$ (truncated at four states) and $C^{\textrm{3pt}}_{\Gamma}(t;\tau)$ (truncated at 2 or 3 states) within a single elimination jackknife process using binned data. [^1] The binning of the data is done in two steps. First, the $O(100)$ measurements made on each configuration are averaged. Second, these configuration averages are further binned using bin sizes ranging between 2–9 resulting in 250–325 data points for the various ensembles. The auto-correlation function calculated using these binned data, which enter in the single elimination jackknife process, show no measurable correlations. Consequently, no augmentation of the single elimination jackknife errors is deemed necessary. All fits are made using the full covariance matrix with respect to the time slices included in the fits. It is, however, restricted to being block diagonal for simultaneous fits to the two-and three-point correlation functions as the off diagonal elements between them are not well-determined with the current statistics.
We also carried out separate fits to the two- and three-point functions and the two sets of results overlap. No anomalous case showing a bias in the simultaneous fits was found. Since the simultaneous fits give more conservative errors and are better motivated as a method, the results are presented using them.
In fits without priors, or wide priors, the $A_{i}$ and $M_{i}$ are predominantly fixed by $C^{\textrm{2pt}}(\tau)$, while the $C^{\textrm{3pt}}_{\Gamma}(t;\tau)$ give $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ including the GSME $\langle 0|O_{\Gamma}^{q}|0\rangle$. In this case the priors are used mainly to stabilize the fits. The drawback, with current statistics, is that there is a whole region in the $\{A_{i},M_{i}\}$ space that gives fits to $C^{\textrm{2pt}}(\tau)$ with similar $\chi^{2}/dof$. Thus, there is considerable range of plausible values for $M_{1}$ with the lowest, theoretically, being the $N(1)\pi(-1)$ or $N(0)\pi(0)\pi(0)$. One way to reduce this uncertainty is to input physically motivated values of the $M_{i}$ through narrow priors.
Ideally, one wants a data driven analysis procedure, i.e., the data and the fit should fix the $M_{i}$. With the current statistics, we are not able to achieve this desired situation. We, therefore, carry out the full analysis using two physically motivated but different values for $M_{1}$:
  1. 1.
    The “standard” fit to $C^{\textrm{2pt}}(\tau)$ and $C^{\textrm{3pt}}(t;\tau)$ employs wide priors for all excited-state amplitudes, $\mathcal{A}_{i}$, and masses, $M_{i}$. In these simultaneous fits, the priors are used only to stabilize the fits. The $A_{0}$ and $M_{i}$, are mainly constrained by $C^{\textrm{2pt}}(\tau)$. The value of $M_{1}$ found in this case is $\gtrsim 1.5$ GeV. This simultaneous fit is labeled $\{4\}$” and $M_{1}$ is loosely identified with the radial excitation $N(1440)$.
  2. 2.
    The $N\pi$” fit uses a narrow prior for $M_{1}$, centered on the non-interacting energy of the lowest allowed, $N(1)\pi(-1)$ or $N(0)\pi(0)\pi(0)$, lattice state on that ensemble. In practice, the energies of these two states are very similar, so they can be considered as one in Eq. (6) with a combined amplitude given by the sum. Throughout this work, for brevity we use the label $N\pi$ for their combined contribution. The priors for the other states are the same as in the “standard” fit. This fit is labeled $\{4^{N\pi}\}$”.
Since the two fit strategies are not distinguished by the $\chi^{2}$, we explain, when quoting the final results, the physics motivations for the choices made. Also, where appropriate, we quote the difference as an estimate of an associated systematic uncertainty. Further details of our implementation of these two fits strategies are given in Refs. [11, 4].
It is important to note that the mass of the $N\pi$ or the $N\pi\pi$ state decreases below that for the first radial excitation $N(1440)$ only for ensembles with $M_{\pi}\lesssim 200$ MeV, reaching $\sim 1230$ MeV for physical pion mass ensembles. Thus, in cases with enhanced matrix elements, one expects to observe a strong dependence on $M_{\pi}$ only for $M_{\pi}<200$ MeV, which in our calculation leaves only the $a09m130$ ensemble. Fits to these data are shown in Figs. 2 and 3.
Refer to caption   Refer to caption     Refer to caption
Figure 2: Data for the bare charges $g_{A,T,S}^{u,d}$ (sum of the connected and disconnected contributions) obtained at various separations $\{\tau,t\}$ are plotted versus $(t-\tau/2)/a$ for the physical $M_{\pi}$ ensemble $a09m130$. We show both the “standard” and $N\pi$” fits, defined in the text, when ESC are manifest. In each case, the result of the fit is shown by lines of the same color as the data for the various $\tau/a$ listed in the label, and the $\tau\to\infty$ value is given by the gray band. We also show, for comparison, the light quark disconnected contributions, $g_{A,S,T}^{l}$, in the panels on the right and the ES fits to them. The $\chi^{2}/dof$ of the fit with the largest AIC score is given within square parenthesis in the top right corner.
Figure 3: Data for the bare strange charges, $g_{A,T,S}^{s}$, obtained at various separations $\{\tau,t\}$, are plotted versus $(t-\tau/2)/a$ for the physical $M_{\pi}$ ensemble $a09m130$. These have only disconnected contributions. Only the “standard” fits to remove ESC are shown and considered since the expected lowest multihadron state, $\Sigma K$, is heavier than the nucleon’s radial excitation. Rest is the same as in Fig. 2.
Furthermore, previous studies indicate that the difference in nucleon isovector axial and tensor charges, $g_{A,T}^{u-d}$, from the two strategies is small [11]. This is also observed for $g_{A,T}^{u,d,s}$ in this study, and the errors in results with the $N\pi$” strategy are comparable to the differences. Only for $g_{S}^{u,d}$ do the differences stand out. Thus, to nail down the best analysis strategy with up to 3-state fits requires future higher statistics studies on more ensembles with $M_{\pi}\lesssim 200$ MeV.
To illustrate the difference in the values of $M_{1}$ coming out of the “standard” and $N\pi$” analyses, we quote the $a09m130$ values, which have the maximum difference. The Bayesian prior and width used for the mass gap $a\Delta M_{1}=a(M_{1}-M_{0})$ is $0.35(20)$ for the “standard” and $0.12(2)$ for the $N\pi$” analysis with the central values corresponding to 0.79 GeV and 0.27 GeV in physical units, respectively. The output $a\Delta M_{1}$ from the simultaneous fits to $C^{\text{2pt}}$ and $C^{\text{3pt}}_{\Gamma}$ are consistent with the input priors, i.e., they lie in the range $0.24\sim 0.30$ for the “standard” and $0.11\sim 0.14$ for the $N\pi$” analysis. We note little dependence of output values on the quark flavor and/or the operator.
IV Model average of various fits made to remove ESC
We have carried out the following twelve ESC fits, i.e., the 12 “models”, for both the “standard” and the $N\pi$” strategies to quantify change under variations of the fit parameters.
  • The full list of $\tau$ values analyzed (covering 0.9–1.5fm) is given in Table 2. The ESC analysis is done (i) with and (ii) without using the data with the largest $\tau$. These largest $\tau$ data have the largest statistical but smallest systematic errors.
  • For each $\tau$, we neglect data on $\mathop{\tau_{\rm skip}}\nolimits$ time slices next to the source and the sink that have the largest ESC. Fits with three different $t_{\text{skip}}$ values in the range of 0.2–0.4 fm were made.
  • The spectral decomposition of the 3-point function, Eq. (5), is truncated at (i) 2-states and (ii) 3-states. In all 3-state fits, the $\langle 2|\mathcal{O}|2\rangle$ term is set to zero as it is not resolved by our data.
For the final results, we take the average weighted by the AIC score times the inverse of the variance, $\propto\text{exp}[-(\chi^{2}-2N_{\text{dof}})/2]/\sigma^{2}$ [13, 37]. In most cases, one set of values of $\tau$ and $t_{\text{skip}}$ gives scores that are much higher than all others for both the 2- and 3-state fit. Thus, in practice, the model average is mainly over the variation of the result with the order of the truncation of Eq. (5). While most 2- and 3-state fits are similar, an example in the $N\pi$” analysis where we note a qualitative difference, even though the AIC scores were similar, is shown in Fig. 4. Since we do not expect an enhanced $N\pi$ state contribution to $g_{A,T}^{q}$, the large extrapolation in the 2-state fit indicates that this is an artifact due to inadequate statistics. The large error and smaller extrapolation in the 3-state fit is further indication that the data, in this and such cases, are statistics limited. More on this point in Sec. V.
Figure 4: The two panels shows the 2-state (top) and 3-state (bottom) $N\pi$” fit to the same data for $g_{A}^{d}$ from the $a09m130$ ensemble. Points with $\tau=\{10,12,14,16\}$ and $t_{\text{skip}}=2$ are included in the fits and are shown using filled circles. These two fits have the largest and similar AIC scores and dominate the model average, but differ qualitatively—the 2-state fit shows a large extrapolation, while the 3-state fit has a large error—and give different central values.
Results from a 2- versus 3-state truncation of Eq. (5) highlight the tension between balancing systematic and statistical errors, which is accentuated in the $N\pi$” analysis. The 2-state fit results may have significant systematic error from residual ESC whereas the statistical precision of the data limits the robustness of the 3-state fits. The same tension is also present between the fits with the two sets of $\tau$ values. The $\chi^{2}/dof$ or the AIC score or the $p$ value of the fit, in most cases, does not distinguish between them. We regard cases with large extrapolation in the ESC fits or a large growth in errors or a large difference between 2- and 3-state fits as outliers due to inadequate statistics rather than a physical effect. The differences are, nevertheless, used to assign an appropriate systematic uncertainty in these cases.
V Pattern of Excited-State Contributions to $g_{A,S,T}^{u,d,s}$
The $a09m130$ ensemble data for the ratio $C^{\text{3pt}}(t;\tau)/C^{\text{2pt}}(\tau)$ are shown in Figs. 2 and  3 as a function of the values of separations $\tau$ and current insertion time $t$ simulated. As already stated above, we choose this physical pion mass ensemble to point out the notable features since the differences in $\Delta M_{1}$ between the “standard” and the $N\pi$” analyses on it are the largest. In each panel, the final result, the AIC weighted average of the twelve ESC fits, the models, is shown by the horizontal grey band. The uncertainty bands on data points with a given color are obtained as follows: The value at each $\{\tau,t\}$ point is the weighted average of the twelve model values using the individual AIC weights, and the ”error” is the spread. These points are then simply joined to form the bands. Note that these bands are not related in any simple way to the final grey band, but are presented to show the variation in the twelve values, which is given by the thickness of the bands.
Theoretically, the data for all $\tau$ must become symmetric about $t-\tau/2$ for the charges as the statistical precision improves, and the convergence should become monotonic for large enough $\tau$. Also, when choosing the $\tau_{\rm max}$ to use in the fits, the errors at that $\tau$ should ideally be small enough to resolve the differences in data at a given $t$ but different $\tau$. Clearly, for the physical mass ensemble $a09m130$, the data are already marginal by these criteria at $\tau/a=14$ ($\approx 1.2$ fm).
In Refs. [11, 2, 3, 4], we had studied the pattern of ESC to $g_{A,S,T}^{u,d}$ separately for the connected and disconnected contributions. Fig. 2 here shows fits to the sum, $C^{\text{conn}}_{\Gamma}(t;\tau)+C^{\text{disc}}_{\Gamma}(t;\tau)$. For completeness, we have compared these results for charges versus those obtained by doing separate fits to the connected and disconnected contributions, and find the differences are smaller than the errors. A summary of our observations is as follows:
  • $g_{A}^{u}$: The convergence of the connected contribution is from below. The disconnected contribution is negative ($\sim-0.06$) and converges to a more negative value as illustrated by $g_{A}^{l,s}$ in Figs. 2 and 3. A notable consequence is that the ESC mostly cancel in the sum and the data at different $\tau$ overlap within the current statistical precision as illustrated for $g_{A}^{u}$ in the top left panel in Fig. 2. We, therefore, estimate the $\tau\to\infty$ value by the weighted average of the data for a range of $\tau$ and $t$ values, i.e., the “plateau method”. Thus, the “standard” and the $N\pi$” results are the same.
  • $g_{A}^{d}$: The connected and disconnected contributions are both negative and their ESC add in contrast to what is observed for $g_{A}^{u}$. We, therefore, perform 2- and 3-state fits to obtain the $\tau\to\infty$ value and the extrapolation is significantly larger in the $N\pi$” analysis.
  • $g_{A}^{l}$ and $g_{A}^{s}$: The disconnected contribution is small and negative. The convergence is towards a more negative value for both as evident from Figs. 2 and 3.
  • $g_{T}^{u}$ and $g_{T}^{d}$: The ESC contribution to $g_{T}^{u}$ ($g_{T}^{d}$) is significant (small). The disconnected contribution is small (<0.01<0.01) and negative. Thus, it only marginally reduces $g_{T}^{u}$ and increases $|g_{T}^{d}|$. More important, the fits to $g_{T}^{u}$ and $g_{T}^{d}$ are good and the size of the extrapolation to the $\tau\to\infty$ value is small.
  • $g_{T}^{l}$ and $g_{T}^{s}$: The values are much smaller than the errors and there is no clear pattern of convergence. We, therefore, assume that there are no significant ESC and make two fits to a constant ansatz—with and without using the full covariance matrix. The results are consistent, however, the fit with the covariance matrix gives a larger error, and is chosen for the final result.
  • $g_{S}^{u}$ and $g_{S}^{d}$: The connected and disconnected contributions to both are positive, the ESC are large, and the convergence is from below. The size of the extrapolation in the $N\pi$” fits is large, giving larger values for $g_{S}^{u,d}$, which implies a larger pion-nucleon sigma term [4].
  • $g_{S}^{l}$ and $g_{S}^{s}$: Their values are also positive and significant, however, the errors in the current data are large. In one case, $g_{S}^{s}$ from the $a09m310$ ensemble, only the diagonal elements of the covariance matrix were used since we did not find reasonable ranges of $\{\tau,t,t_{\rm skip}\}$ for which the off-diagonal elements were stable.
Results for the bare charges extracted using the two strategies to remove ESC, and after model averaging, are summarized in Tables 35 and 4. Their renormalization and conversion to the $\overline{\text{MS}}$}} scheme at the scale $2$ GeV is done using the analysis for the 2+1-flavor theory described in Appendix B.
[…]
The error in the $\tau\to\infty$ extrapolated value for $g_{A}^{u,d}$ in the $N\pi$” analysis is large. For $g_{A}^{u}$, we see no significant ESC, so the “standard” fit value is taken. In $g_{A}^{d}$, the ESC to the connected and disconnected terms add and the $N\pi$” fit value is about $30\%$ larger. We do not have a theoretical reason for such a large effect, nevertheless we take the average of the two estimates. Note, results for the isovector axial charge $g_{A}^{u-d}$, which has no disconnected contributions, are consistent with our final values already given in Ref. [12] as shown in Table 9.
[…]
The $g_{A,T}^{s}$, illustrated in Fig. 3, get only disconnected contributions and the lowest multihadron state is likely to be $\Sigma K$ as illustrated in Fig. 1). This is heavier than the radial excitation of the nucleon. Also, the data are noisy and the values are small. For these reasons we choose the “standard” analysis. It also gives an error estimate that we consider conservative.
[…]
Lastly, even though we have chosen the “standard” analysis values for $g_{A,S}^{s}$, we do show the CCFV extrapolation of the data from the $N\pi$” analysis in Figs. 10 and 16. The procedure for the renormalization of these charges is discussed in Appendix B.
Ensemble $g_{A}^{u-d}$ $g_{A}^{u}$ $g_{A}^{d}$ $g_{A}^{u+d}$ $g_{A}^{l}$ $g_{A}^{s}$
a15m310 $\{3\}$ 1.2416(57) 0.892(12) -0.3538(57) 0.535(13) -0.0431(28) -0.0277(16)
$\{3^{N\pi}\}$ 1.2507(82) 0.892(12) -0.3646(76) 0.516(16) -0.0510(39) -0.0315(22)
a12m310 $\{4\}$ 1.267(11) 0.8896(97) -0.372(24) 0.469(29) -0.0696(77) -0.0401(48)
$\{4^{N\pi}\}$ 1.2741(79) 0.8896(97) -0.372(21) 0.479(20) -0.0699(64) -0.0420(27)
a12m220 $\{4\}$ 1.285(30) 0.886(17) -0.415(28) 0.457(43) -0.083(16) -0.049(13)
$\{4^{N\pi}\}$ 1.304(23) 0.886(17) -0.441(28) 0.425(55) -0.098(21) -0.0595(72)
a09m310 $\{4\}$ 1.256(45) 0.8695(87) -0.3614(78) 0.506(14) -0.0532(57) -0.0304(25)
$\{4^{N\pi}\}$ 1.251(11) 0.8695(87) -0.3688(85) 0.497(16) -0.0578(66) -0.0336(35)
a09m220 $\{4\}$ 1.307(80) 0.8558(93) -0.414(17) 0.441(24) -0.084(15) -0.0448(60)
$\{4^{N\pi}\}$ 1.318(33) 0.8558(93) -0.436(25) 0.382(30) -0.115(17) -0.0620(85)
a09m130 $\{4\}$ 1.281(32) 0.842(13) -0.441(32) 0.386(46) -0.100(24) -0.047(13)
$\{4^{N\pi}\}$ 1.406(42) 0.842(13) -0.618(48) 0.104(77) -0.227(35) -0.118(34)
a06m310 $\{4\}$ 1.196(45) 0.870(31) -0.372(23) 0.488(47) -0.067(20) -0.0396(49)
$\{4^{N\pi}\}$ 1.227(31) 0.862(29) -0.379(18) 0.468(37) -0.073(13) -0.0429(55)
a06m220 $\{4\}$ 1.256(32) 0.842(24) -0.390(17) 0.434(34) -0.079(15) -0.0476(70)
$\{4^{N\pi}\}$ 1.271(32) 0.843(28) -0.396(19) 0.426(38) -0.094(26) -0.0515(85)
Table 3: Final bare axial charges $g_{A}^{u,d,s}$. These $g_{A}^{u-d}$ results are obtained by doing the ES fit to the difference of the connected data with the insertion of the $u$ and $d$ quarks to take into account the correlations between the two. The $g_{A}^{u}$ and $g_{A}^{d}$ are extracted from a fit to the sum of the connected plus disconnected correlation functions. Their sum, $g_{A}^{u}+g_{A}^{d}$, can be compared to the result of a fit to the sum of the correlation functions, $g_{A}^{u+d}$. The errors are calculated using a single elimination jackknife method using data binned as described in the text, and each number is the AIC weighted average over the various model values. The results from the “standard” and $N\pi$” ESC fits are given separately.
Ensemble $g_{T}^{u-d}$ $g_{T}^{u}$ $g_{T}^{d}$ $g_{T}^{u+d}$ $g_{T}^{l}$ $g_{T}^{s}$
a15m310 $\{3\}$ 1.1231(59) 0.902(12) -0.2278(39) 0.669(11) -0.0075(16) -0.00204(54)
$\{3^{N\pi}\}$ 1.1096(80) 0.889(14) -0.2211(47) 0.654(14)
a12m310 $\{4\}$ 1.0608(81) 0.827(25) -0.2221(57) 0.601(34) -0.00517(94) -0.00188(55)
$\{4^{N\pi}\}$ 1.0509(96) 0.837(18) -0.2219(52) 0.612(25)
a12m220 $\{4\}$ 1.062(13) 0.865(21) -0.2087(92) 0.657(25) -0.0046(23) -0.0020(10)
$\{4^{N\pi}\}$ 1.039(28) 0.846(25) -0.198(13) 0.648(36)
a09m310 $\{4\}$ 1.0288(47) 0.8210(84) -0.2086(34) 0.6099(83) -0.0056(10) -0.00159(63)
$\{4^{N\pi}\}$ 1.033(19) 0.822(15) -0.2080(45) 0.6030(95)
a09m220 $\{4\}$ 0.997(13) 0.791(15) -0.2015(40) 0.568(26) -0.0047(19) -0.00178(69)
$\{4^{N\pi}\}$ 0.982(20) 0.773(21) -0.1941(46) 0.538(39)
a09m130 $\{4\}$ 1.003(11) 0.765(25) -0.2114(96) 0.528(29) -0.0113(32) -0.0017(15)
$\{4^{N\pi}\}$ 1.038(86) 0.741(74) -0.164(25) 0.48(11)
a06m310 $\{4\}$ 0.985(15) 0.8005(68) -0.2023(38) 0.5907(81) -0.0057(14) -0.00201(74)
$\{4^{N\pi}\}$ 0.962(31) 0.788(22) -0.201(12) 0.581(26)
a06m220 $\{4\}$ 0.961(18) 0.757(20) -0.186(10) 0.562(21) -0.0058(16) -0.00155(81)
$\{4^{N\pi}\}$ 0.933(34) 0.756(27) -0.182(12) 0.557(27)
Table 4: Final bare tensor charges $g_{T}^{u,d,s}$. The rest is the same as in Table 3.
Ensemble $g_{S}^{u-d}$ $g_{S}^{u}$ $g_{S}^{d}$ $g_{S}^{u+d}$ $g_{S}^{l}$ $g_{S}^{s}$
a15m310 $\{3\}$ 0.829(24) 3.83(11) 3.068(87) 6.91(20) 0.648(77) 0.701(36)
$\{3^{N\pi}\}$ 0.829(31) 3.94(12) 3.26(12) 7.17(23) 0.761(70) 0.709(32)
a12m310 $\{4\}$ 0.906(24) 4.64(23) 3.95(33) 8.54(54) 1.10(24) 0.694(80)
$\{4^{N\pi}\}$ 0.893(58) 4.57(15) 3.73(14) 8.28(28) 0.978(72) 0.671(47)
a12m220 $\{4\}$ 1.037(88) 5.70(31) 4.53(22) 10.23(50) 1.15(23) 0.664(70)
$\{4^{N\pi}\}$ 1.13(11) 6.22(33) 4.98(28) 11.33(57) 1.53(25) 0.719(64)
a09m310 $\{4\}$ 1.008(22) 4.47(14) 3.77(28) 8.10(34) 0.73(18) 0.499(57)
$\{4^{N\pi}\}$ 1.047(33) 4.73(21) 3.81(21) 8.55(43) 0.88(15) 0.597(82)
a09m220 $\{4\}$ 1.008(31) 5.69(42) 4.68(43) 10.34(83) 1.19(38) 0.69(13)
$\{4^{N\pi}\}$ 1.001(62) 6.24(41) 5.18(41) 11.38(81) 1.53(30) 0.87(15)
a09m130 $\{4\}$ 0.89(10) 6.37(50) 5.70(56) 12.0(1.0) 1.30(27) 0.71(14)
$\{4^{N\pi}\}$ 0.95(27) 9.17(81) 8.74(88) 17.5(1.5) 3.25(75) 1.11(30)
a06m310 $\{4\}$ 1.180(70) 5.78(54) 4.53(43) 10.47(98) 0.99(15) 0.64(10)
$\{4^{N\pi}\}$ 1.258(89) 5.78(49) 4.59(36) 10.47(84) 1.12(16) 0.69(12)
a06m220 $\{4\}$ 0.818(88) 6.15(61) 6.2(1.2) 12.8(1.7) 1.51(31) 0.520(96)
$\{4^{N\pi}\}$ 0.83(11) 6.19(57) 6.01(90) 12.7(1.4) 1.67(38) 0.62(20)
Table 5: Final bare scalar charges $g_{S}^{u,d,s}$. The rest is the same as in Table 3.
VI Chiral-Continuum-Finite-volume (CCFV) extrapolation and error budget
The renormalized axial, $g_{A}^{u,d,s}$ and tensor, $g_{T}^{u,d,s}$ charges are extrapolated to the physical point, $a\to 0$, $M_{\pi}=135$ MeV, and $M_{\pi}L\to\infty$, using the ansatz
$g(a,M_{\pi},M_{\pi}L)=c_{0}+c_{a}a+c_{2}M_{\pi}^{2}+c_{3}\frac{M_{\pi}^{2}e^{-M_{\pi}L}}{\sqrt{M_{\pi}L}}\,$, (7)
that includes the leading corrections, pertinent to our lattice setup, in all three variables $\{a,M_{\pi},M_{\pi}L\}$.
For the scalar charges $g_{S}^{u}$ and $g_{S}^{d}$, chiral perturbation theory gives two differences. First, the chiral behavior, $g_{S}^{u,d}=d_{0}+d_{a}a+d_{1}M_{\pi}+d_{2}M_{\pi}^{2}+M_{\pi}^{2}\log M_{\pi}^{2}+\ldots$, starts with a term proportional to $M_{\pi}$ [39]. Second, the contribution of $N\pi$ and $N\pi\pi$ excited states from higher order terms (chiral logs, etc.) is large as discussed in [4]. The difference on including these in ESC fits is shown in Fig. 2. Thus, one needs to include the higher order terms (chiral logs, etc.) in both, fits to remove possibly large ESC and in the chiral part of the CCFV ansatz. Unfortunately, with data at only three values of $M_{\pi}$, only the first two terms, $\propto M_{\pi}$ and $M_{\pi}^{2}$ in addition to lattice spacing dependence, can be kept without overparameterization. Noting these points, the CCFV ansatz used for the scalar charges $g_{S}^{u,d}$ is
$g_{S}^{u,d}(a,M_{\pi},M_{\pi}L)$ $=d_{0}+d_{a}a+d_{2}M_{\pi}+d_{3}M_{\pi}^{2}$
$+d_{4}M_{\pi}\left(1-\frac{2}{M_{\pi}L}\right)e^{-M_{\pi}L}\,.$ (8)
Note that the finite-volume term is also modified [4]. For the strange scalar charge $g_{S}^{s}$, and for $g_{\{A,T\}}^{s}$, we do not expect a large contribution from pion loops since the lightest multihadron state is expected to be the much heavier $\Sigma K$. So we use the ansatz given in Eq. (7).
The final results for the charges, after CCFV (or CC) extrapolations for the two strategies and the two renormalization methods, are summarized in Table 6. From these we obtain the final central values and the error budget given in Table 7. Adding errors in quadrature, the final results are given in Table 8. We show, in Figs. 816 in Appendix A, only the CCFV fits but point out that the finite volume correction is small in all cases as can be inferred from the values given in Table 6. In panels showing CCFV fits versus $M_{\pi}^{2}$, the additional grey bands show fits keeping only the chiral corrections given in Eqs. (7) and (8). Overlap between these grey bands and the CCFV fits (pink bands), implies that the other two systematics, discretization and finite volume corrections, are small. Overall, no significant finite-volume corrections are observed for $M_{\pi}L\gtrsim 4$.
Some specific observations on these CCFV fits to obtain the charges at the physical point are as follows.
VI.1 Axial
CCFV fits in Fig. 8 and 9 for $g_{A}^{u}$ and $g_{A}^{d}$, respectively, show similar large positive slope versus $M_{\pi}^{2}$, but with a larger uncertainty in the $N\pi$” analysis. There also is a significant positive slope in $g_{A}^{u}$ versus $a$ but not in $g_{A}^{d}$. These variations decrease $g_{A}^{u}$ and make $g_{A}^{d}$ more negative as $a\to 0$ and $M_{\pi}\to 135$ MeV. They add in $g_{A}^{u+d}$, making it smaller, and almost cancel in $g_{A}^{u-d}$.
The CCFV estimates for $g_{A}^{u}$, summarized in Table 6, are, within errors, independent of the ESC strategy and the renormalization method. We therefore, take the weighted average of the four values and the largest error to get $g_{A}^{u}=0.781(22)$. To this we add a systematic uncertainty of 0.011 to account for the difference between the CCFV and CC values and quote $g_{A}^{u}=0.781(22)(11)_{\rm CC}$ in Table 7.
Estimates of $g_{A}^{d}$ are consistent between the two renormalization methods but $\approx 10\%$ more negative with the $N\pi$” analysis. Based on the reasoning given in the extraction of the isovector axial form factors in [11, 12], the $N\pi$” analysis is suggested, however, given the size of the errors, we defer choosing. So we take the mean and quote half the difference, $0.024$, as a systematic uncertainty along with $0.009$ for the difference between the the CCFV and CC values. This gives $g_{A}^{d}=-0.440(29)(9)_{\rm CC}(24)_{N\pi}$.
For $g_{A}^{s}$, we choose the “standard” analysis as discussed previously and assign a systematic uncertainty only for the CCFV versus CC fits to get $g_{A}^{s}=0.055(9)(1)_{\rm CC}$. These results for $g_{A}^{u,d,s}$, summarized in Table 7 and 8, are consistent with, and update those published in Ref. [2].
VI.2 Tensor
The magnitude of ESC in $g_{T}^{u}$ is significant, while the errors in $g_{T}^{d}$ and $g_{T}^{s}$ are large, especially on the physical pion mass ensemble $a09m130$, as shown in Fig. 2 and in Ref. [3]. The CCFV fits are shown in Fig. 11, 12 and 13. The “standard” and $N\pi$” analyses give consistent estimates but the errors in the $N\pi$” case are $\approx 50\%$ larger. For all three charges, $g_{T}^{u,d,s}$, we average the values from the two renormalization methods and use the difference between CCFV and CC fits to assign a systematic uncertainty. Since no enhanced contributions from multihadron, $N\pi$”, states is expected in the tensor channel, our choice is the “standard” analysis and the CCFV fits. Nevertheless, to be conservative and account for the $\approx 2\sigma$ difference between the “standard” and $N\pi$ values for $g_{T}^{d}$, we assign half the difference, 0.012, as an additional systematic uncertainty. The final results are summarized in Tables 7 and 8.
VI.3 Scalar
[…]
VI.4 The pion-nucleon sigma term $\sigma_{\pi N}$
[…]
$\sigma_{\pi N}=(d_{2}+d_{2}^{a}a)M_{\pi}^{2}+d_{3}M_{\pi}^{3}+d_{4}M_{\pi}^{4}+d_{4L}M_{\pi}^{4}\log\frac{M_{\pi}^{2}}{M_{N}^{2}}\,.$ (9)
[…]
In this work, we also do the CCFV extrapolation of the renormalized charges $g_{S}^{u+d}$ and $g_{S}^{u,d,s}$. Our preferred value for the isoscalar charge, $g_{S}^{u+d}$, is obtained by doing ESC fits to the sum of the total $u$ and $d$ correlators. The results are $g_{S}^{u+d}=12.4(1.2)$ for the “standard” and $g_{S}^{u+d}=17.8(1.6)$ for the $N\pi$” analysis. The value $g_{S}^{s}=0.37(14)$ is, as explained before, taken from the standard analysis. Combining these with the renormalized quark masses $m_{l}=3.427(51)$ and $m_{s}=93.46(58)$ MeV given in the FLAG 2024 report [5] for the $N_{f}=2+1+1$ theory, provides the second estimates: $\sigma_{\pi N}|_{N\pi}=61(6)$ MeV for the $N\pi$ analysis, and $\sigma_{\pi N}|_{\rm standard}=43(4)$ MeV and $\sigma_{s}=35(13)$ MeV from the “standard” analysis. Note that if the individual values of $g_{S}^{u}$ and $g_{S}^{d}$ are taken from Table 7, then $\sigma_{\pi N}$ is larger by about 2 MeV.
Averaging the two estimates for $\sigma_{\pi N}$ and taking the larger of the errors, our final values are
$\sigma_{\pi N}|_{\rm standard}$ $=42(6)~{\rm MeV}\,$,
$\sigma_{\pi N}|_{N\pi}$ $=61(6)~{\rm MeV}\,$,
$\sigma_{s}|_{\rm standard}$ $=35(13)~{\rm MeV}\,.$ (10)
These are summarized in Fig. 7 along with other lattice determinations with 2+1+1 and 2+1-flavor simulations that meet the FLAG criteria [5], and also the results for $\sigma_{\pi N}$ from phenomenology.
“Standard” analysis for removing ESC $N\pi$” analysis
$q$ $g_{A}^{q}$ $g_{T}^{q}$ $g_{S}^{q}$ $g_{A}^{q}$ $g_{T}^{q}$ $g_{S}^{q}$
Extrapolation excluding finite volume corrections (CC fit). $\rm{Z_{1}}$ method
u 0.803(17) 0.784(19) 6.35(56) 0.804(18) 0.776(28) 9.28(84)
d -0.430(17) -0.2001(65) 6.03(62) -0.468(21) -0.1903(87) 8.79(89)
s -0.0539(64) -0.00164(92) 0.457(95) -0.0725(78) -0.00165(92) 0.73(12)
Extrapolation excluding finite volume corrections (CC fit). $\rm{Z_{2}}$ method
u 0.779(18) 0.759(19) 6.32(57) 0.779(18) 0.747(29) 9.37(87)
d -0.419(17) -0.1939(68) 6.03(63) -0.458(22) -0.1830(91) 8.91(91)
s -0.0543(69) -0.0016(10) 0.41(10) -0.0738(85) -0.0016(10) 0.68(13)
Extrapolation including finite volume corrections (CCFV fit). $\rm{Z_{1}}$ method
u 0.784(22) 0.787(25) 6.36(56) 0.785(22) 0.776(37) 9.34(85)
d -0.417(27) -0.1957(93) 6.05(62) -0.465(29) -0.172(13) 8.81(91)
s -0.0540(88) -0.0016(11) 0.38(12) -0.0705(95) -0.0016(11) 0.66(16)
Extrapolation including finite volume corrections (CCFV fit). $\rm{Z_{2}}$ method
u 0.776(22) 0.778(26) 6.32(57) 0.778(22) 0.768(38) 9.38(88)
d -0.414(27) -0.1949(96) 6.04(63) -0.462(29) -0.171(13) 8.87(93)
s -0.0554(94) -0.0015(12) 0.36(13) -0.073(10) -0.0015(12) 0.66(17)
Table 6: Results for the flavor diagonal charges obtained after extrapolation to the physical point with and without finite volume corrections (the CCFV and CC fits). Results are presented for the two strategies (“standard” and $N\pi$”) for removing ESC and the two methods ($\rm{Z_{1}}$ and $\rm{Z_{2}}$, defined in Appendix B.4) for determining the renormalization constants. Note the tiny difference in $g_{T}^{s}$ between the “standard” and $N\pi$” analysis since the mixing from light flavors depends on $N\pi$.
This work PNDME 2018
$q$ $\phantom{-}g_{A}^{q}$ $\phantom{-}g_{T}^{q}$ $g_{S}^{q}|_{N\pi}$ $g_{S}^{q}|_{\rm St.}$ $\phantom{-}g_{A}^{q}$ [2] $\phantom{-}g_{T}^{q}$ [3]
$u$ $\phantom{-}0.781(22)(11)_{\rm CC}$ $\phantom{-}0.782(26)(11)_{\rm CC}$ $9.36(88)(4)_{\rm CC}$ $6.34(57)(1)_{\rm CC}$ $\phantom{-}0.777(25)(30)$ $\phantom{-}0.784(28)(31)$
$d$ $0.440(29)(9)_{\rm CC}(24)_{N\pi}$ $0.195(10)(2)_{\rm CC}(12)_{N\pi}$ $8.84(93)(1)_{\rm CC}$ $6.04(63)(1)_{\rm CC}$ 0.438(18)(30)-0.438(18)(30) 0.204(11)(10)-0.204(11)(10)
$s$ $055(9)(1)_{\rm CC} 0.0$$0016(12)(1)_{\rm CC}$ $0.66(17)(5)_{\rm CC}$ $37(13)(6)_{\rm CC} 0.0$53(8)-0.053(8) 0.027(16)-0.027(16)
Table 7: Final results for the flavor diagonal charges obtained by combining the values given in Table 6 using criteria specified in the text. The subscript $C$ denotes the systematic uncertainty assigned using the difference between the CCFV and CC fit values; and $N\pi$ for that between the “standard” and $N\pi$” strategies where appropriate. Values with the $\rm{Z_{1}}$ and $\rm{Z_{2}}$ renormalization methods (defined in Appendix B.4) are averaged. Results for $g_{A}^{q}$ and $g_{T}^{q}$ supersede those published previously in Refs. [2] and [3] and reproduced here.
Final results
$q$ $\phantom{-}g_{A}^{q}$ $\phantom{-}g_{T}^{q}$ $g_{S}^{q}$
$u$ $\phantom{-}0.781(25)$ $\phantom{-}0.782(28)$ $9.39(88)$
$d$ 0.440(39)-0.440(39) 0.195(16)-0.195(16) $8.84(93)$
$s$ 0.055(9)-0.055(9) 0.0016(12)-0.0016(12) $0.37(14)$
Table 8: Final results for the flavor diagonal charges obtained by combining, in quadrature, the errors given in Table 7. The results for the axial charges, $g_{A}^{u,d}$, are the average over the $N\pi$” and “standard” analysis; the tensor, $g_{T}^{u,d}$, and the strange charges, $g_{A,S,T}^{s}$, are from the “standard” analysis; and $g_{S}^{u,d}$ are from the $N\pi$” analysis as explained in the text.
VI.5 Extracting the isovector charges $g_{A,T,S}^{u-d}$
The disconnected contributions, which have large errors, cancel exactly in the isovector charges, $g_{A,T,S}^{u-d}$, for the isospin-symmetric theory. Thus, our overall strategy to get the isovector charges is to make ES fits to the difference of the connected contributions, i.e., to $C^{\text{u,conn}}_{\Gamma}(t;\tau)-C^{\text{d,conn}}_{\Gamma}(t;\tau)$. These fits are much more robust compared to those for the individual flavor diagonal charges, and benefit from correlated fluctuations that are missed in the construction $g^{u,bare}-g^{d,bare}$. For example, extracting $g_{S}^{u-d}$ from separate fits to the two $C^{\text{conn+disc}}_{S}(t;\tau)$ gives $O(1)$ errors with current data. This occurs for the two reasons illustrated by the $a09m130$ data in Fig. 5. First, the data for the larger $\tau$ ($\tau=14$ and $16$) do not satisfy the two conditions that must hold for fits to give consistent results—monotonic convergence from below and symmetry about $\tau/2$. These features are exhibited by the data from the other seven ensembles and we have checked that they improve with increasing statistical precision. Second, within errors, which increase rapidly with $\tau$, the data around $\tau/2$ overlap for $\tau=8,10,12$. This, ideally, indicates convergence by the plateau criteria, however, for confirmation much higher precision $\tau=14$ and $16$ data are needed. On the other hand, using 2-state fits to remove residual ESC not well-constrained.
Figure 5: The ratio plot for $g_{S}^{u-d,\text{bare}}$ on the a09m130 ensemble. The $\tau=\infty$ value is taken from the fit to $\tau=8,10$ and 12 data, which show no significant ESC at $t=\tau/2$. The $\tau=14$ and 16 data are neglected as, with current statistics, they have large errors and deviate from the expected behavior of monotonic convergence.
Our final results for isovector charges from the full set of clover-on-HISQ calculations (using 13 data points on 11 HISQ ensembles, CCFV extrapolation, and taking the average of values with $\rm{Z_{1}}$ and $\rm{Z_{2}}$ renormalization) have already been published in Ref. [12]. For completeness, Table 9 summarizes the values from this work using eight ensembles, “standard” and $N\pi$” analyses, CCFV and CC extrapolations; and the same two renormalization methods. The two sets of results are consistent, with the values obtained here having slightly larger errors. The only notable feature in both is the difference between the various estimates for $g_{A}^{u-d}$ that has already been pointed out in Ref. [12].
“Standard” analysis for removing ESC $N\pi$” analysis
This work This work This work Ref. [12] This work This work This work Ref. [12]
($\rm{Z_{1}}$) ($\rm{Z_{2}}$) ($\rm{Z_{1}}$) ($\rm{Z_{2}}$)
$g_{A}^{u-d}|_{\rm CC}$ 1.284(29) 1.221(30) 1.328(26) 1.277(27)
$g_{A}^{u-d}|_{\rm CCFV}$ 1.221(39) 1.206(40) 1.214(40) 1.240(27) 1.288(32) 1.278(33) 1.283(33) 1.294(49)
$g_{T}^{u-d}|_{\rm CC}$ 1.007(12) 0.969(12) 0.976(27) 0.929(28)
$g_{T}^{u-d}|_{\rm CCFV}$ 1.010(16) 0.998(17) 1.004(17) 0.989(34) 0.976(36) 0.969(37) 0.973(37) 0.991(24)
$g_{S}^{u-d}|_{\rm CC}$ 0.99(11) 0.98(11) 1.17(27) 1.17(27)
$g_{S}^{u-d}|_{\rm CCFV}$ 0.98(11) 0.96(12) 0.97(12) 1.022(100) 1.18(27) 1.17(28) 1.18(28) 1.085(115)
Table 9: Results for the isovector charges with the two strategies, “standard” and $N\pi$”, used to remove ESC; the two methods, $\rm{Z_{1}}$ and $\rm{Z_{2}}$, used for renormalization; and the two ansatz, CCFV and $C$, used to extrapolate data to the physical point. The final values from this work are compared with the average of $\rm{Z_{1}}$ and $\rm{Z_{2}}$ values from Ref. [12] obtained with CCFV extrapolation of the full set of thirteen clover-on-HISQ calculations. The CCFV extrapolations based on eight versus thirteen points account for most of the differences between the results. The “standard” analysis strategy is labeled $3^{\ast}$” in Ref. [12], while the $N\pi$” analysis is similar to “3-RD”.
[…]
VII Comparison with Previous Results and Conclusions
Our final results, which supersede those presented in Refs. [2] and [3], are summarized in Tables 7 and  8. The improvements made in this work are described in the introduction I. The FLAG 2024 report [5] quotes our previous results for $g_{A,T}^{u,d,s}$ (from Refs. [2, 3]) as the FLAG values for 2+1+1-flavor simulations. The updated versions of the FLAG summary figures, including data from other collaborations that meet the FLAG criteria for averaging, and post FLAG 2024 results from the ETM collaboration [40], are shown in Figs. 6 and 7.
From the axial charges, $g_{A}^{u,d,s}$ given in Table 8, we calculate the contribution of the intrinsic spin of quarks to the nucleon spins to be $0.5\Delta\Sigma_{q}=0.5\sum_{q}g_{A}^{q}=0.143(24)$. This is consistent with the COMPASS analysis $0.13<0.5\Delta\Sigma_{q}<0.18$ evaluated at 3 GeV2 [57].
The lattice determination of tensor charges, $g_{T}^{u,d,s}$, is far more accurate than extractions from the transversity distributions of quarks in semi-inclusive deep-inelastic scattering (SIDIS) experiments at Jefferson Lab and other facilities worldwide [58, 59, 60]. These charges also give the intrinsic contributions of the electric dipole moments of the quarks to the neutron/proton EDM [3].
[…]
The most direct comparison of our results is with the latest results from the ETM collaboration 24 [40], who have analyzed three physical mass ensembles using 2+1+1-flavors of twisted mass fermions. We note that these new ETM results supersede all their previously published values as explained in [40].
The two sets of results, summarized in Figs. 6 and 7, are consistent with two notable ($2\sigma$) differences: in $g_{A}^{l,s}$ and $\sigma_{\pi N}$. The reasons are:
  • Our data show ESC in $g_{A}^{l,s}$ (see Figs. 2 and 3) while the ETM data (labelled $g_{A}^{u+d(disc)}$ and $g_{A}^{s(disc)}$) are flat even though they include data at smaller $\tau$, i.e., in the range $0.41<\tau<1.6$ fm versus $0.9<\tau<1.5$ fm in our calculation. This accounts for most of the difference in the final values of $g_{A}^{u,d}$ shown in Fig. 6.
[…]
In summary, as already remarked in Sec. V, the largest uncertainty in these two quantities in our work comes from the ESC fits and possible contributions from $N\pi$ states. The ETM results are from a “standard” analysis.
Looking ahead, we are analyzing two physical pion mass HISQ ensembles at $a\approx 0.087$ and $0.057$ fm with about six times larger statistics (6000 and 4000 configurations, respectively) and better tuned values of $m_{s}$ and $m_{l}$ in the clover action. These calculations will improve control over excited-state analysis and, hopefully, the statistics are large enough to provide a data driven resolution between the “standard” and $N\pi$” analyses.
Acknowledgements.
We thank the MILC collaboration for providing the 2+1+1-flavor HISQ lattices, and acknowledge John Gracey for discussions on the renormalization of the currents. The calculations used the Chroma software suite [26]. This research used resources at (i) the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; (ii) the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725, through the ALCC program project LGT107 and HEP145, and the INCITE program project HEP133; (iii) the USQCD collaboration, which is funded by the Office of Science of the U.S. Department of Energy; and (iv) Institutional Computing at Los Alamos National Laboratory. This work was prepared in part by LLNL under Contract DE-AC52-07NA27344. S. Park acknowledges the support from the ASC COSMON project. S. Park acknowledges support from the U.S. DOE Contract No. DE-AC05-06OR23177, under which Jefferson Science Associates, LLC, manages and operates Jefferson Lab. Also acknowledged is support from the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. DOE Office of Science and the National Nuclear Security Administration. T. Bhattacharya and R. Gupta were partly supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Contract No. DE-AC52-06NA25396. T. Bhattacharya, R. Gupta, S. Mondal, S. Park, and B. Yoon were partly supported by the LANL LDRD program, and S. Park by the Center for Nonlinear Studies.
Appendix A Chiral-continuum-finite-volume (CCFV) extrapolation for $g_{A,S,T}^{u,d,s}$
Figures 816 in this appendix show the CCFV fits used to extrapolate the data for the charges on the eight ensembles to the physical point. In each column, the three panels show the simultaneous CCFV fit (pink band) plotted versus either $a$ or $M_{\pi}^{2}$ or $M_{\pi}L$, respectively, with the other two variables set to their physical point values. In the panels plotted versus $M_{\pi}^{2}$, the additional grey bands show the fit with only the chiral corrections included in Eqs. (7) or (8). Note that when the discretization and finite volume corrections are small, the grey bands overlap with the CCFV fits (pink bands). The results of these fits are summarized in Table 6.
[…]
Figure 9: CCFV extrapolation for $g_{A}^{d}$ using Eq. (7). The rest is the same as in Fig. 8.
Figure 10: CCFV extrapolation for $g_{A}^{s}$ using Eq. (7). The rest is the same as in Fig. 8.
Figure 11: CCFV extrapolation for $g_{T}^{u}$ using Eq. (7). The rest is the same as in Fig. 8
Figure 12: CCFV extrapolation for $g_{T}^{d}$ using Eq. (7). The rest is the same as in Fig. 8
Figure 13: CCFV extrapolation for $g_{T}^{s}$ using Eq. (7). The rest is the same as in Fig. 8. These data are only analyzed using the “standard” strategy as the likely multihadron lowest excited state, $\Sigma K$, is heavier than the $N(1440)$ radial excitation.
Figure 14: CCFV extrapolation for $g_{S}^{u}$ using Eq. (8). The rest is the same as in Fig. 8.
Figure 15: CCFV extrapolation for $g_{S}^{d}$ using Eq. (8). The rest is the same as in Fig. 8.
Figure 16: CCFV extrapolation for $g_{S}^{s}$. The rest is the same as in Fig. 8
[…]
Appendix B Nonperturbative renormalization for the 2+1-flavor theory using the RI-sMOM scheme
This Appendix presents an overview of the calculation of the renormalization constants for the flavor diagonal quark bilinear operators $\mathcal{O}^{f}=\bar{\psi}^{f}\Gamma\psi^{f}$ with Dirac matrix $\Gamma$ and the flavor index $f=\{u,d,s\}$ in the $N_{f}=3$ theory. The lattice calculation is done in the RI-sMOM scheme [8, 9] on four ensembles at the different values of the lattice spacing used in this study: $a15m310$, $a12m310$, $a09m310$, and $a06m310$. Since data are obtained at only one value of $M_{\pi}\neq 0$ at each $a$, we explicitly assume that the quark mass dependence can be neglected and these results are a good approximation to the chiral limit values. Second, the lattice calculation is not fully $O(a)$ improved in neither the action nor the operators. In the latter, we neglect all the lattice spacing dependent improvement terms. Consequently, the results for the renormalized charges are extrapolated to the continuum limit using an ansatz starting with a term linear in $a$.
Results for the matching factors between the RI-sMOM and the $\overline{\text{MS}}$} scheme, and the running in the $\overline{\text{MS}}$} scheme are taken from continuum perturbation theory [62, 63, 64, 65, 66, 67] as specified in Appendix B.3.
The general relation between renormalized, $\mathcal{O}_{R}$, and bare, $\mathcal{O}$, operators including mixing between flavors is:
$\mathcal{O}^{f}_{R}$ $=\sum_{f^{\prime}}Z_{\Gamma}^{ff^{\prime}}\mathcal{O}^{f^{\prime}}\,$, (11)
with the matrix $Z_{\Gamma}^{ff^{\prime}}$ determined nonperturbatively on the lattice using the regularization independent (RI) renormalization scheme [8] in which the renormalized vertex function is set to its tree-level value. The calculation is done with the gauge fields fixed to the Landau gauge and quark propagators generated using momentum sources. The final results for the three charges, using the two methods, labeled $\rm{Z_{1}}$ and $\rm{Z_{2}}$ and discussed in Appendix B.4, for partially removing discretization errors are given in Appendix B.5.
B.1 Flavor mixing in the RI scheme
The calculation starts with the amputated vertex function $\Gamma^{ff^{\prime}}(p_{1},p_{2})$ defined as,
$\Gamma^{ff^{\prime}}(p_{1},p_{2})$ $=\langle S^{f}(p_{1})\rangle^{-1}\langle\psi^{f}(p_{1})\mathcal{O}^{f^{\prime}}\psi^{f}(p_{2})\rangle\langle S^{f}(p_{2})\rangle^{-1}$ (12)
where $\psi^{f}$ and $S^{f}$ are the quark field and propagator with flavor-$f$. The three-point function $\langle\psi^{f}(p_{1})\mathcal{O}^{f^{\prime}}\psi^{f}(p_{2})\rangle$ gets contributions from both the connected and the disconnected diagrams shown in Fig. 18. The perturbative result for this Green’s function $\Gamma^{ff^{\prime}}(p_{1},p_{2})$ in the $\overline{\text{MS}}$} scheme is given in Refs. [68, 64] in terms of its Lorentz’s components.
[…]
$Tr[\mathbb{P}_{\Gamma}\Gamma^{ff^{\prime}}_{R}(p_{1},p_{2})]\equiv\frac{Z_{\Gamma}^{f^{\prime}f^{\prime\prime}}}{Z_{\psi}^{f}}Tr[\mathbb{P}_{\Gamma}\Gamma^{ff^{\prime\prime}}(p_{1},p_{2})]=\delta^{ff^{\prime}}\,.$ (13)
This relation defines the $3\times 3$ mixing matrix $Z_{\Gamma}$ for the 3-flavor theory in the RI-sMOM scheme:
$Z_{\Gamma}=\left[\frac{1}{Z_{\psi}^{f}}Tr({\mathbb{P}_{\Gamma}\Gamma(p_{1},p_{2})}^{T})\right]^{-1}\,.$ (14)
The determination of $Z_{\psi}^{f}$ is done in a separate calculation using two different methods as discussed in Sec. B.4.
(a) The connected vertex $\Gamma_{\text{conn}}(m,\mu^{2})$
(b) The disconnected vertex $\Gamma_{\text{disc}}(m_{f},m_{f^{\prime}},\mu^{2})$
Figure 18: The amputated vertex diagrams defined in Eq. (12)
The amputated vertex function $\Gamma^{ff^{\prime}}$ gets contributions from both the connected and disconnected diagrams shown in Fig. 18:
$\Gamma^{ff^{\prime}}(p_{1},p_{2})=\Gamma_{\text{conn}}(m^{f},\mu^{2})\delta^{ff^{\prime}}-\Gamma_{\text{disc}}(m^{f},m^{f^{\prime}},\mu^{2}),$ (15)
The (-) sign in $\Gamma_{\text{disc}}$ is due to the anticommuting nature of the fermion fields contracted into the quark loop. On the lattice we calculate the projected amputated vertex functions $c_{f}$ (connected) and $d_{ff^{\prime}}$ (disconnected) defined as
$c_{\Gamma}^{f}$ $=\frac{1}{Z_{\psi}^{f}}\text{Tr}\left[\mathbb{P}_{\Gamma}\Gamma_{\text{conn}}(m^{f},\mu^{2})\right]$ (16)
$d_{\Gamma}^{ff^{\prime}}$ $=\frac{1}{Z_{\psi}^{f}}\text{Tr}\left[\mathbb{P}_{\Gamma}\Gamma_{\text{disc}}(m_{f},m_{f^{\prime}},\mu^{2})\right]\,.$ (17)
For the $N_{f}=2+1$ isospin symmetric theory relevant to this work, the determination of $Z_{\Gamma}^{ff^{\prime}}$ require calculating the following 6 quantities on the lattice,
$c_{\Gamma}^{l},~c_{\Gamma}^{s},~d_{\Gamma}^{ll},~d_{\Gamma}^{ls},~d_{\Gamma}^{sl}\text{, and }d_{\Gamma}^{ss}$ (18)
for each $\Gamma$. Working in the flavor basis $f\in\{u-d,u+d,s\}$, the $Z_{\Gamma}$, defined in Eq. (14), is block diagonal:
$Z_{\Gamma}$ $\equiv\begin{pmatrix}Z_{\Gamma}^{u-d,u-d}&0&0\\ 0&Z_{\Gamma}^{u+d,u+d}&Z_{\Gamma}^{u+d,s}\\ 0&Z_{\Gamma}^{s,u+d}&Z_{\Gamma}^{ss}\\ \end{pmatrix}$ (19)
$=\begin{pmatrix}c_{\Gamma}^{l}&0&0\\ 0&c_{\Gamma}^{l}-2d_{\Gamma}^{ll}&-2d_{\Gamma}^{sl}\\ 0&-d_{\Gamma}^{ls}&c_{\Gamma}^{s}-d_{\Gamma}^{ss}\\ \end{pmatrix}^{-1}.$ (20)
For the discussion of the matching and running, the more convenient basis is the isovector (3), octet (8), and singlet (0) one, i.e., $f\in\{u-d,u+d-2s,u+d+s\}$. In this case the renormalization matrix in terms of the 6 quantities is
$\hskip-22.76228pt\left(\begin{array}[]{ccc}Z^{3,3}_{\Gamma}&0&0\\ 0&Z^{8,8}_{\Gamma}&Z^{8,0}_{\Gamma}\\ 0&Z^{0,8}_{\Gamma}&Z^{0,0}_{\Gamma}\end{array}\right)=\left(\begin{array}[]{ccc}c_{\Gamma}^{l}&0&0\\ 0&\frac{1}{3}(c_{\Gamma}^{l}+2(c_{\Gamma}^{s}-d_{\Gamma}^{ll}+d_{\Gamma}^{ls}+d_{\Gamma}^{sl}-d_{\Gamma}^{ss}))&\frac{2}{3}(c_{\Gamma}^{l}-c_{\Gamma}^{s}-2d_{\Gamma}^{ll}+2d_{\Gamma}^{ls}-d_{\Gamma}^{sl}+d_{\Gamma}^{ss})\\ 0&\frac{1}{3}(c_{\Gamma}^{l}-c_{\Gamma}^{s}-2d_{\Gamma}^{ll}-d_{\Gamma}^{ls}+2d_{\Gamma}^{sl}+d_{\Gamma}^{ss})&\frac{1}{3}(2c_{\Gamma}^{l}+c_{\Gamma}^{s}-4d_{\Gamma}^{ll}-2d_{\Gamma}^{ls}-2d_{\Gamma}^{sl}-d_{\Gamma}^{ss})\end{array}\right)^{-1}\,.$ (27)
B.2 RI-sMOM scheme
The lattice calculation is done in the RI-sMOM scheme [9] where the 4-momentum of the external legs $\{p_{1},p_{2}\}$ satisfies the symmetric momentum condition $p_{1}^{2}=p_{2}^{2}=(p_{1}-p_{2})^{2}=\mu^{2}$, and $\mu^{2}$ defines the renormalization scale. We find that the matrix $Z^{\text{RI-sMOM}}(\mu)$ is close to diagonal with $c_{\Gamma}^{l}\sim O(1)$ and $d_{\Gamma}^{ff^{\prime}}$ about a few percent at $\mu\sim 2$ GeV. This is illustrated in Fig. 19 for the scalar operator (largest mixing) for 3 of the disconnected projected amputated Green’s functions $d_{\Gamma}^{lf}$ calculated at various $\mu$. As expected, the value decreases as the quark mass in the loop is increased from light to strange to charm. To get a signal for this small mixing, we use momentum source propagators [69] in Landau gauge and take the momenta $\{p_{1},p_{2}\}$ to be proportional to $(1,1,1,1)$ to minimize $O(4)$-symmetry breaking. A comparison of the improvement in the extraction of $Z$ using momentum versus point sources is shown in Fig 20.
Figure 19: Data for the disconnected projected amputated Green’s function $d_{\Gamma}^{ff^{\prime}}$ for a quark with flavor $f\in\{l,s,c\}$ and the disconnected loop with flavor $f^{\prime}$ for the scalar bilinear operator $O_{S}^{f^{\prime}}$ in the RI-sMOM scheme. The allowed momenta satisfy $\mu^{2}=p^{2}={p^{\prime}}^{2}=(p-p^{\prime})^{2}$. Data for the charm quark are also plotted to show the dependence on the quark mass even though they are not used in this work.
Figure 20: These plots show the improvement in the calculation of the four renormalization factors in the RI-sMOM scheme obtained using momentum (mom src) versus point source propagators (point src). The data are for the a06m310 ensemble. For comparison, we also show the factors obtained in the RI-MOM scheme even though they are not used in this work.
B.3 Matching and RG running
The calculation of the renormalization factors for each operator involves the four steps described below.
  1. 1.
    To obtain the full $3\times 3$ matrix, $Z^{\text{RI-sMOM,{latt}}}(\mu)$ defined in Eq. (20), the two $c_{f}$ and four $d_{ff^{\prime}}$ (Fig. 19) are calculated using momentum source propagators with the lattice fixed to Landau gauge. The simulations are done for a range of lattice momenta satisfying $q^{2}=p_{1}^{2}=p_{2}^{2}\equiv\mu^{2}_{\rm latt}$.
  2. 2.
    For each $\mu_{\rm latt}$, a horizontal matching, $\mu_{\rm cont}=\mu_{\rm latt}$, to $\overline{\text{MS}}$} scheme is performed by multiplying $Z^{\text{RI-sMOM,{latt}}}(\mu)$ with the matching factor $C^{\text{RI-sMOM}\to{\overline{\text{MS}}}}(\mu)$ calculated using perturbation theory in the continuum:
    $\hskip 14.45377ptZ^{{\overline{\text{MS}}}}(\mu)=C^{\text{RI-sMOM}\to{\overline{\text{MS}}}}(\mu)Z^{\text{RI-sMOM,{latt}}}(\mu)\,.$ (28)
  3. 3.
    The result for each $\mu$ is run in the continuum to a common scale, which we choose to be 2 GeV:
    $\hskip 14.45377ptZ^{{\overline{\text{MS}}}}(2\mathop{\rm GeV}\nolimits;\mu)=C^{{\overline{\text{MS}}}}(2\mathop{\rm GeV}\nolimits,\mu)Z^{{\overline{\text{MS}}}}(\mu)\,.$ (29)
  4. 4.
    Lattice artifacts giving rise to a $\mu$ dependence. These are removed using the ansatz
    $\hskip 14.45377ptZ^{{\overline{\text{MS}}}}(2\mathop{\rm GeV}\nolimits;\mu)=Z^{{\overline{\text{MS}}}}(2\mathop{\rm GeV}\nolimits)+c_{1}\mu^{2}+c_{2}\mu^{4}.$ (30)
    Note that, in the RI-sMOM calculation, there are, in general, not enough data points at small $\mu^{2}$ to include a $1/\mu^{2}$ term as was done in Refs. [70, 19]. We, therefore, fit to points with $\mu^{2}_{\rm min}$ large enough so that the contribution of a $1/\mu^{2}$ term can be assumed to be negligible.
The final results for $Z^{{\overline{\text{MS}}}}_{\{A,S,T\}}(2\mathop{\rm GeV}\nolimits)$ are given in Appendix B.5. We remind the reader of the implicit assumption that the possible quark mass dependence has been neglected and the results obtained at $M_{\pi}\neq 0$ are a good approximation to the chiral limit values.
In the continuum calculation of $C^{\text{RI-sMOM}\to{\overline{\text{MS}}}}(\mu)$ and $C^{{\overline{\text{MS}}}}(2\mathop{\rm GeV}\nolimits,\mu)$ for the scalar, tensor and pseudoscalar operators, the singlet and non-singlet factors are the same because the disconnected loops vanish in the zero mass limit due to chirality. Thus the $3\times 3$ matrices for the matching and running reduce to an overall factor, for example, $C^{\text{RI-sMOM}\to{\overline{\text{MS}}}}(\mu)=Z^{\rm match}\times\mathds{1}$. The three-loop result for the scalar and tensor can be obtained, after projection, from the amputated Green’s functions given in Ref. [62]:
$\hskip-17.07182ptZ^{match}_{S,\text{SMOM}}$ $=$ $1+0.64552a\ +\ (23.024\,-4.0135n_{f})a^{2}$
$+$ $\left(2.1844n_{f}^{2}-169.923n_{f}+889.74\right)a^{3},$
$\hskip-17.07182ptZ^{match}_{T,\text{SMOM}}$ $=$ $1-0.215a\ +\ (4.103n_{f}-43.384)a^{2}$ (31)
- $\left(7.0636n_{f}^{2}-309.828n_{f}+1950.76\right)a^{3},$
where $a(\mu)\equiv\alpha_{s}(\mu)/4\pi$. The running factor, $C^{{\overline{\text{MS}}}}(2\mathop{\rm GeV}\nolimits,\mu)$ for the tensor is obtained from the four-loop results in Ref. [64], and for the scalar the five-loop result in Ref. [63] .
The matching and running factors are unity for both the singlet and non-singlet vector operator because of the conservation of the vector current. The renormalization factor $Z_{V}$ for the bare local operator is given by the relation $Z_{V}g_{V}=1$, with $g_{V}$ obtained from the forward matrix element as discussed in Appendix B.4.
For the axial operator, only the flavor nonsinglet axial current is conserved in the chiral limit. Therefore the matching from RI-sMOM to $\overline{\text{MS}}$} scheme and the RG running factor is unity for the nonsinglet operator.
The flavor singlet axial current is not conserved due to the $U_{A}(1)$ anomaly, and the renormalization factors become nontrivial at the two-loop level. This $1+\mathcal{O}(\alpha^{2})$ correction is taken from Refs. [71, 68], and the three-loop anomalous dimension from Ref. [72] is used for the RG running. [^2] The 4-loop anomalous dimension for the isosinglet axial current in the continuum in the $\overline{\text{MS}}$} scheme has been calculated in Ref. [73]. We use the 3-loop running for consistency with the RI-sMOM to $\overline{\text{MS}}$} matching that has been calculated to only 2-loops.
The off-diagonal terms in the mixing matrix for the singlet axial current are zero in the chiral limit, leaving only the diagonal elements. The difference between the singlet and nonsinglet (the disconnected contribution) has been calculated in the $\overline{\text{MS}}$} scheme in the RI-sMOM setup in Ref. [68]. This is used to get the following matching factor for the axial current in the $\{3,8,0\}$ basis:
$\hskip-19.91684ptC_{A}^{\text{RI-sMOM}\to{\overline{\text{MS}}}}(\mu)=\left(\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1-26.5\,a^{2}\,n_{f}\end{array}\right)$ (35)
In the calculation of all the matching and running factors, the number of flavors used is $n_{f}=4$. We have checked that the difference in the final results between using $n_{f}=3$ and $4$ is below the claimed precision.
On the lattice, Wilson-clover fermions break chiral symmetry explicitly and there are non-zero contributions from disconnected loops for all charges. These give rise to off-diagonal terms as shown in Eqs. (20) and (27). The mixing in the tensor channel is $O(\alpha_{s}^{3})$ in lattice perturbation theory [74] implying $r_{T}\equiv Z_{T}^{\text{singlet}}/Z_{T}^{\text{non-singlet}}=1+O(\alpha_{s}^{3})\approx 1$ where $a=\alpha_{s}/4\pi$ and $n_{f}$ denotes the number of flavors. The largest disconnected contribution (off-diagonal elements $d_{\Gamma}^{ff^{\prime}}$) is for the axial. In the $\{u-d,u+d,s\}$ basis, it is approximately $7\%$ for the coarsest $a\approx 0.15$ fm ensemble and decreases to below 2% at the finest $a\approx 0.06$ fm ensemble.
The final factors used to renormalize the charges are given in Appendix B.5. Note the significant differences between results with the $\rm{Z_{1}}$ versus the $\rm{Z_{2}}$ strategies (defined in Appendix B.4) used to calculate $Z_{\psi}$. There is a similar pattern of differences between the three diagonal elements with $Z^{u-d,u-d}\approx Z^{u+d,u+d}$ that differ significantly from $Z^{s,s}$. This mass dependence is discussed in Appendix B.4 and our current understanding is, it is a discretization effect that in the final analysis is, to a large extent, taken care of in the continuum (CCFV) extrapolation of the renormalized charges. The mixing is $O(\alpha_{s}^{2})$ in lattice perturbation theory [74], which would be larger than what the lattice data for the tensor operator given in Appendix B.5 suggests.
In Fig. 21, we show the data for $Z_{\Gamma}$ in $\overline{\text{MS}}$} scheme at $2\mathop{\rm GeV}\nolimits$ for the $a06m310$ ensemble as a functions of $\mu$ along with fits using Eq. (30). The results for the diagonal parts, for both $\rm{Z_{1}}$ and $Z_{2}$, show a difference that vanishes in the continuum limit. The off-diagonal mixing elements $Z_{\Gamma}^{u+d,s}$ and $Z_{\Gamma}^{s,u+d}$ shown in the bottom two rows are all smaller than $1\%$ and results with the $\rm{Z_{1}}$ and $\rm{Z_{2}}$ strategies essentially overlap.
One advantage of using RI-sMOM scheme for the scalar channel is the better convergence between 2- and 3-loop running and smaller deviation from unity in the perturbative series for the matching factor compared to the RI-MOM scheme [66, 9] as illustrated in Fig. 22.
To summarize, for our setup, the numerical results given in Appendix B.5 show that the off-diagonal terms and the $r_{\cal O}-1$ factors are small.
Figure 21: Data for $Z_{\Gamma}$ in $\overline{\text{MS}}$} scheme at $2\mathop{\rm GeV}\nolimits$ on the a06m310 ensemble along with a quadratic fit versus $\mu^{2}$ defined in Eq. (30). Each panel compares results from the two different strategies, $\rm{Z_{1}}$ (blue) and $\rm{Z_{2}}$ (red). The four columns give results for the axial, scalar, tensor and vector operators, respectively. Each row represents one of the 5 nonzero matrix elements in the $3\times 3$ matrix defining $Z$ in the $\{u-d,u+d,s\}$ basis in Eq. (20).
Figure 22: The top left panel shows the matching factor $C^{\text{RI-sMOM}\to{\overline{\text{MS}}}}=Z^{\text{RI-sMOM}}(\mu)/Z^{{\overline{\text{MS}}}}(\mu)$ for the singlet axial current. It acquires a non-trivial dependence on the scale $\mu$ starting at 2-loops. For scalar and tensor operators, we compare the 2- and 3-loop matching factors for RI-SMOM $\to{\overline{\text{MS}}}$ and RI-MOM $\to{\overline{\text{MS}}}$. Bottom Right: The 4-loop running of $\alpha/4\pi$ versus $\mu$ for the $N_{f}=4$ theory used in this study.
B.4 Two renormalization strategies $\rm{Z_{1}}$ and $\rm{Z_{2}}$ defined by the choice of $Z_{\psi}^{q}$
To define the $\rm{Z_{1}}$ and $\rm{Z_{2}}$ renormalization strategies used, recall that the renormalization constants for the isovector bilinear operators in the RI scheme are given by
$Z_{\Gamma}(p)=\frac{Z_{\psi}(p)}{c_{\Gamma}(p)}\quad{\rm or}\quad Z_{\psi}(p)=c_{\Gamma}(p)Z_{\Gamma}(p)$ (36)
where $p$ is the momentum insertion, $Z_{\psi}$ is the renormalization constant for the fermion field, and $c_{\Gamma}$ is the projected amputated connected 3-point function calculated in Landau gauge as defined in Eq. (16). $\rm{Z_{1}}$ and $\rm{Z_{2}}$ are defined by the two ways used to calculate $Z_{\psi}$.
  • $\rm Z_{1}$: The $Z_{\psi}$ is calculated from the projected bare quark propagator
    $Z_{\psi}(p)=\frac{i}{12p^{2}}\operatorname{Tr}[S_{B}^{-1}(p)\text{$\not\!p$}]$ (37)
  • $\rm Z_{2}$ calculated using the vector Ward identity (VWI), $Z_{V}g_{V}=1$. Equation (11) then becomes
    $Z_{\psi}^{\rm VWI}(p)=c_{V}(p)/g_{V},$ (38)
    where $c_{V}$ is the projected amputated connected 3-point function with insertion of the local vector operator within the quark state while the vector charge $g_{V}$ is obtained from the insertion of the same operator within any, in principle, hadronic state, i.e., pion, nucleon, etc. We use the nucleon state.
These two strategies were used in Refs. [70, 19] for isovector bilinear operators of light quarks, i.e., $g_{\Gamma}Z_{\Gamma}$ and $Z_{\Gamma}/Z_{V}\times g_{\Gamma}/g_{V}$. Later, they were called $\rm Z_{1}$ and $\rm Z_{2}$ in Ref. [11]. In Refs. [70, 19] we showed that they have different behavior versus $q^{2}$ as the discretization effects are different in $Z_{\psi}$ and $Z_{\psi}^{\rm VWI}$, which are obtained from $S_{B}(p)$ and $c_{\Gamma}^{V}(p)$, respectively [^3] The $O(a)$ improvement for $S_{B}(p)$ is studied in Ref. [75], and $c_{\Gamma}^{V}(p)$ (off-shell improvement) is studied in Ref. [76]. To within expected discretization errors, we checked
  • $Z_{V}|_{\rm Z_{1}}\times g_{V}^{l,{\rm bare}}=1$ for $m_{l}$ as shown in Table 10. The deviation from unity decreases as $a\to 0$. When this relation is satisfied, renormalization using $\rm Z_{1}$ and $\rm{Z_{2}}$ should give consistent results.
  • Final continuum limit values of all the nucleon charges (and form factors), using $\rm Z_{1}$ and $\rm Z_{2}$ agree within the quoted errors [11].
This work extends the calculation to the heavier strange quark. The data in Tab. 10 show that the deviation of the renormalized $g_{V}^{s}$ from unity is significantly larger. While this deviation for both the light and strange quarks goes to zero essentially linearly as $a\to 0$ as shown in Figs. 23 and 24, it is important to note that the discretization artifact is large and has a significant mass dependence.
We point out that these deviations are not due to incomplete removal of ESC in the extraction of the isovector vector charge $g_{V}$. The data for the bare vector charge, i.e., the ratio plots versus $\{\tau,t\}$, are shown in Fig. 25. The ESC are less than a percent, while the deviations in the data, shown in Tab. 10, are a 2–8% effect for even the light quarks.
Figure 23: Results for $Z_{V}$ using the two strategies $\rm{Z_{1}}$ and $\rm{Z_{2}}$.
Figure 24: Data for $g_{V}^{f}Z_{V}^{f}|_{\rm Z_{1}}$, with the vector renormalization factor $Z^{f}_{V}|_{\rm Z_{1}}=1/c^{f}_{V}$, for flavor $f\in\{l,s\}$. See Eq. (36) for details. The isovector vector charge $g_{V}^{\{l,s\}}$ is obtained from the forward matrix element within the nucleon ground state constructed using light and strange valence quark propagators respectively. The trend in the data is consistent with the vector Ward identity (VWI) relation $g_{V}Z_{V}|_{\rm Z_{1}}=1$ being satisfied in the continuum limit.
Ensemble $g_{V}^{l,\text{bare}}$ $Z_{V}^{l}|_{\rm Z_{1}}$ $g_{V}^{l}|_{\rm Z_{1}}$ $g_{V}^{s,\text{bare}}$ $Z_{V}^{s}|_{\rm Z_{1}}$ $g_{V}^{s}|_{\rm Z_{1}}$
a15m310 1.0689(19) 0.86503(28) 0.925(2) 0.9845(3) 0.85690(18) 0.8436(3)
a12m220 1.0630(73) 0.90611(6) 0.963(7) 1.0007(3) 0.90048(4) 0.9011(3)
a09m220 1.0498(15) 0.92538(4) 0.971(1) 1.0053(1) 0.91608(3) 0.9209(1)
a06m310 1.0402(26) 0.94550(6) 0.984(2) 1.0116(3) 0.94110(4) 0.9520(3)
Table 10: The bare and the $\rm Z_{1}$-renormalized vector charge for the light and strange quarks on the four ensembles. The non-perturbative calculations was done using the RI-sMOM scheme. The factor $Z_{V}^{\{l,s\}}|_{Z_{1}}$ is obtained using the quadratic extrapolation ansatz given in Eq. (30). The bare isovector vector charge $g_{V}^{\{l,s\},{\rm bare}}$ is obtained from the forward matrix element within the nucleon ground state.
Figure 25: Data for the nucleon vector charge $g_{V}$ on four HISQ ensembles, $a15m310$, $a12m220$, $a09m220$, and $a06m310$, plotted from left to right. The top row is for the light valence quark with mass $m_{l}$ and the bottom for the strange quark with mass $m_{s}$. The data at the largest three source-sink separations $\tau$ agree at the sub-percent level. For the final result (gray band), we take the unweighted average over 4–10 data points centered about $t-\tau/2=0$ with the largest two $\tau$. Residual ES effects are neglected.
B.5 Final results for the renormalization factors
The final renormalization factors in the $\overline{\text{MS}}$} scheme at 2 GeV, calculated using the 2+1 theory in the flavor basis $f\in\{u-d,u+d,s\}$, are given below for the four ensembles, a15m310, a12m310, a09m310, and a06m310, the three charges, and the two renormalization methods, $Z_{1}$ and $Z_{2}$. These factors are the aggregate of the four steps described in Sec. B.3 and used to renormalize the bare charges for all the ensembles at a given lattice spacing $a$ assuming their dependence on the quark masses is negligible compared to the other uncertainties.
$a$ $\approx 0.15\mathop{\rm fm}\nolimits$
$Z_{A}|_{\rm Z_{1}}$ $=\begin{pmatrix}0.9467(1)&0&0\\ 0&0.8905(21)&-0.0560(27)\\ 0&-0.0269(11)&0.9125(13)\\ \end{pmatrix}$ $Z_{A}|_{\rm Z_{2}}$ $=\begin{pmatrix}1.0222(3)&0&0\\ 0&0.9627(22)&-0.0647(28)\\ 0&-0.0285(12)&1.0803(14)\\ \end{pmatrix}$ (39)
$Z_{S}|_{\rm Z_{1}}$ $=\begin{pmatrix}1.0034(2)&0&0\\ 0&0.9642(75)&-0.0323(75)\\ 0&-0.0198(39)&0.9806(38)\\ \end{pmatrix}$ $Z_{S}|_{\rm Z_{2}}$ $=\begin{pmatrix}1.0792(3)&0&0\\ 0&1.0381(81)&-0.0370(79)\\ 0&-0.0208(42)&1.1554(40)\\ \end{pmatrix}$ (40)
$Z_{T}|_{\rm Z_{1}}$ $=\begin{pmatrix}0.8925(1)&0&0\\ 0&0.8930(6)&0.0003(8)\\ 0&0.0001(3)&0.8764(4)\\ \end{pmatrix}$ $Z_{T}|_{\rm Z_{2}}$ $=\begin{pmatrix}0.9653(2)&0&0\\ 0&0.9659(6)&0.0004(8)\\ 0&0.0001(3)&1.0393(4)\\ \end{pmatrix}$ (41)
$a$ $\approx 0.12\mathop{\rm fm}\nolimits$
$Z_{A}|_{\rm Z_{1}}$ $=\begin{pmatrix}0.95705(4)&0&0\\ 0&0.9287(6)&-0.0283(11)\\ 0&-0.0141(4)&0.9386(4)\\ \end{pmatrix}$ $Z_{A}|_{\rm Z_{2}}$ $=\begin{pmatrix}0.9933(1)&0&0\\ 0&0.9642(6)&-0.0307(11)\\ 0&-0.0143(4)&1.0389(4)\\ \end{pmatrix}$ (42)
$Z_{S}|_{\rm Z_{1}}$ $=\begin{pmatrix}0.9183(1)&0&0\\ 0&0.9163(43)&-0.0074(63)\\ 0&-0.0010(24)&0.9148(30)\\ \end{pmatrix}$ $Z_{S}|_{\rm Z_{2}}$ $=\begin{pmatrix}0.94423(4)&0&0\\ 0&0.9493(44)&-0.0080(64)\\ 0&-0.0010(25)&1.0119(31)\\ \end{pmatrix}$ (43)
$Z_{T}|_{\rm Z_{1}}$ $=\begin{pmatrix}0.94254(5)&0&0\\ 0&0.9424(3)&0.0003(5)\\ 0&-0.0002(2)&0.9358(2)\\ \end{pmatrix}$ $Z_{T}|_{\rm Z_{2}}$ $=\begin{pmatrix}0.97970(4)&0&0\\ 0&0.9792(3)&0.0003(5)\\ 0&-0.0002(2)&1.0380(2)\\ \end{pmatrix}$ (44)
$a$ $\approx 0.09\mathop{\rm fm}\nolimits$
$Z_{A}|_{\rm Z_{1}}$ $=\begin{pmatrix}0.96407(4)&0&0\\ 0&0.9356(17)&-0.0286(18)\\ 0&-0.0155(9)&0.9454(9)\\ \end{pmatrix}$ $Z_{A}|_{\rm Z_{2}}$ $=\begin{pmatrix}0.99252(4)&0&0\\ 0&0.9631(17)&-0.0307(18)\\ 0&-0.0159(9)&1.0251(10)\\ \end{pmatrix}$ (45)
$Z_{S}|_{\rm Z_{1}}$ $=\begin{pmatrix}0.92167(5)&0&0\\ 0&0.9158(40)&-0.0067(40)\\ 0&-0.0025(21)&0.9141(21)\\ \end{pmatrix}$ $Z_{S}|_{\rm Z_{2}}$ $=\begin{pmatrix}0.9464(1)&0&0\\ 0&0.9419(42)&-0.0073(41)\\ 0&-0.0025(22)&0.9900(21)\\ \end{pmatrix}$ (46)
$Z_{T}|_{\rm Z_{1}}$ $=\begin{pmatrix}0.98950(4)&0&0\\ 0&0.9892(6)&-0.0002(7)\\ 0&0.0002(3)&0.9763(4)\\ \end{pmatrix}$ $Z_{T}|_{\rm Z_{2}}$ $=\begin{pmatrix}1.01847(4)&0&0\\ 0&1.0182(6)&-0.0002(7)\\ 0&0.0002(3)&1.0592(4)\\ \end{pmatrix}$ (47)
$a$ $\approx 0.06\mathop{\rm fm}\nolimits$
$Z_{A}|_{\rm Z_{1}}$ $=\begin{pmatrix}0.96958(5)&0&0\\ 0&0.9493(14)&-0.0184(15)\\ 0&-0.0096(9)&0.9579(9)\\ \end{pmatrix}$ $Z_{A}|_{\rm Z_{2}}$ $=\begin{pmatrix}0.98501(4)&0&0\\ 0&0.9652(14)&-0.0192(15)\\ 0&-0.0097(9)&1.0068(9)\\ \end{pmatrix}$ (48)
$Z_{S}|_{\rm Z_{1}}$ $=\begin{pmatrix}0.86298(5)&0&0\\ 0&0.8601(25)&-0.0043(27)\\ 0&-0.0027(13)&0.8606(13)\\ \end{pmatrix}$ $Z_{S}|_{\rm Z_{2}}$ $=\begin{pmatrix}0.8785(1)&0&0\\ 0&0.8743(26)&-0.0045(27)\\ 0&-0.0028(13)&0.9042(13)\\ \end{pmatrix}$ (49)
$Z_{T}|_{\rm Z_{1}}$ $=\begin{pmatrix}1.03069(5)&0&0\\ 0&1.0350(3)&0.0004(3)\\ 0&-0.0000(2)&1.0266(2)\\ \end{pmatrix}$ $Z_{T}|_{\rm Z_{2}}$ $=\begin{pmatrix}1.0501(1)&0&0\\ 0&1.0521(3)&0.0004(3)\\ 0&-0.0000(2)&1.0790(2)\\ \end{pmatrix}$ (50)